### v3

# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies relying on mathematical models can struggle with real-world disruptions, leading to a significant deviation from their anticipated efficiency. This work applies the cutting-edge concept of antifragility to design a traffic control strategy for urban road networks against disruptions. Antifragility sets itself apart from robustness and resilience as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile terms composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network. Promising results demonstrate (a) the superior performance of the proposed algorithm under both static and incremental magnitude of disruptions compared to the state-of-the-art algorithms, and (b) under incremental disruptions, the proposed method exhibits less fragility compared to the baseline methods. The proposed antifragile methodology has the potential to be integrated into systems subject to disruptions across various disciplines.
keywords:
antifragility , reinforcement learning , perimeter control , disruptions , macroscopic fundamental diagram
journal: Elsevier
[…]
1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations (Auer et al., 2016). Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness and resilience of transportation systems is highly crucial (Ganin et al., 2019). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Dickerson et al., 2000). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with a linearly increasing magnitude of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and exponentially growing negative effects.
To address such issues, the concept of antifragility has shed light on a potential solution. First introduced in the reputable general publication Antifragile: Things That Gain from Disorder (Taleb, 2012), and later mathematically formulated as academic works in Taleb (2013); Taleb and Douady (2013), it has provided insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated earlier in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks. Ever since the proposal of antifragility, it has gained substantial interest from both the public and academia, particularly in the complex systems (Axenie et al., 2024) and risk engineering communities (Aven, 2015; Grassi et al., 2024). However, the design principles for realizing antifragile transportation systems are generally unexplored. One promising approach to induce antifragility is applying learning-based algorithms, such as Reinforcement Learning (RL) (Haydari and Yılmaz, 2022), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
The main goal of this paper is to design an antifragile perimeter control algorithm against different types of traffic disruptions with various magnitudes. The key contributions are: a) We formulate and distinguish the concept of antifragility as opposed to other related and commonly used terms in the transportation domain; b) We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods; and c) We propose a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under environment with increasing disruptions.
The remainder of this paper is structured as follows. In Section 2, we introduce relevant literature on multiple aspects covered in this work, while Section 3 mathematically formulates the cordon-shaped perimeter control simulation environment. Methodologies related to incorporating antifragility into RL algorithms are detailed in Section 4. Section 5 discusses the simulation setups and parametrization. Results are presented in Section 6, followed by conclusion remarks and further discussions in 7.
2 Literature review
This section reviews the relevant literature with three topics intersecting in this work. First, a macroscopic traffic model and the control strategy applied in this paper are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we study how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, the definitions of robustness, resilience, and antifragile are distinguished while introducing the design philosophy of an antifragile system.
2.1 Macroscopic fundamental diagram and perimeter control
Alleviating urban network congestion can be realized through various traffic control strategies. Since Geroliminis and Daganzo (2008) demonstrated the presence of the Macroscopic Fundamental Diagram (MFD) with empirical data and Daganzo and Geroliminis (2008) generated the first MFD analytically with variational theory for a homogeneous urban area, the mathematical relationship between flow-density-speed has paved the path for the development of control strategies at a macroscopic level, enabling more computationally feasible real-time traffic control strategies for large-scale networks (Knoop et al., 2012), such as perimeter control (Keyvan-Ekbatani et al., 2012; Geroliminis et al., 2013; Kouvelas et al., 2017; Yang et al., 2017), congestion pricing (Zheng et al., 2012; Zheng and Geroliminis, 2016; Genser and Kouvelas, 2022), route guidance (Yildirimoglu et al., 2015; Fu et al., 2022).
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free. One major issue with the previous works is the assumption of MFD homogeneity. To tackle this challenge, a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for a sub-network (Ambühl et al., 2019; Saedi et al., 2020), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) and Ji et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers.
2.2 Leveraging RL to induce robustness or resilience
Given the presence of real-world disturbances and disruptions, it is essential to assess the robustness and resilience of any newly developed traffic control algorithm. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged these features to enhance the robustness and resilience of RL-based traffic control strategies. Here, we present a collection of studies that apply RL in traffic control and claim to demonstrate robustness or resilience. While Zhou and Gayah (2021); Chen et al. (2022); Su et al. (2023); Zhou and Gayah (2023) focus on perimeter control frameworks, Aslani et al. (2018); Rodrigues and Azevedo (2019); Chu et al. (2020); Tan et al. (2020); Wu et al. (2020); Korecki et al. (2023) explore the design for signalized intersections. Some of these works directly compare the performance of the proposed algorithm with the benchmark methods under disruptions to showcase either robustness or resilience, whereas Rodrigues and Azevedo (2019); Tan et al. (2020); Chu et al. (2020); Zhou and Gayah (2023) further state explicitly how it is induced with modifications of the algorithm. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are given additional information related to disturbances or disruptions of the environment. The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evokes unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately since they might indicate a shift in the environment. In this case, exploration refers to the process of looking for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness and resilience can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between robustness, resilience, and antifragility. We follow the definitions proposed by Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. In Taleb (2012), two levels of antifragility have been indicated: proto-antifragility and antifragility. To differentiate these two terms, we rename the latter and the more important one as progressive antifragility. Proto-antifragility describes systems that can improve performance when exposed to moderate disruptions with magnitudes in a similar range, resembling the concept of adaptiveness. An illustrated example can be the biological process of hormesis. The higher level of antifragility, i.e., progressive antifragility, emphasizes the concave response of the system under an increasing magnitude of disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is shown in Fig. 13. Although Sun et al. (2024) has proven mathematically that urban road networks are progressively fragile by nature, it should be noted that they can be less fragile when proper control strategies are adopted to induce antifragility.
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines, such as energy (Coppitters and Contino, 2023), biology (Kim et al., 2020), medicine (Axenie et al., 2022), and robotics (Axenie and Saveriano, 2023). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Taleb and West, 2023; Axenie et al., 2022). With this additional information, the system can anticipate ongoing disruptions and be more responsive to drastic changes. Similar to the function of redundancy in resilience (Tan et al., 2019; Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility de Bruijn et al. (2020); Johnson and Gheorghe (2013); Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, current studies on antifragility in engineering mostly pertain to proto-antifragility, such as post-disaster treatment and reconstruction efforts, as in Fang and Sansavini (2017); Priyadarshini et al. (2022). Leveraging the potential of antifragility for the operation of transportation systems is still a largely novel and unexplored notion.
3 Problem Formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 1(a). Additional traffic demand at time $t$ for an Origin-Destination (OD) pair from region $i$ to region $j$ is denoted as $q_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference to accommodate vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 1(b). Given the total number of vehicles with presence in region $i$ at time $t$, denoted as $n_{i}(t)$, the total trip completion rate for this region $i$, denoted as $M_{i}(t)$, can be determined using the corresponding MFD, which comprises both the intraregional trip completion, i.e., $M_{ii}(t)$ and the interregional transfer flow, i.e., $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. To protect regions from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to go across the perimeter at time $t$ is regulated by two perimeter controllers denoted as $u_{ij}(t)\>(i\neq j)$. A list of all notations used in this paper, including the notations used in defining the RL algorithm and the antifragile terms, is summarized in Table 1.
(a) The cordon-shaped urban network
(b) MFDs for the inner and outer regions.
Figure 1: The network structure and the related MFDs.
Table 1: List of notations
Symbol Description
1. General notations in problem formulation
$t$ Time
$\Delta t$ Time step interval
$T$ Total simulation time
$n_{ij}(t)$ Vehicle accumulation with OD from region $i$ to $j$ at time $t$
$n_{i}(t)$ Vehicle accumulation in region $i$ at time $t$
$u_{ij}(t)$ Perimeter control variables regulating flow from region $i$ to $j$ at time $t$
$q_{ij}(t)$ Traffic demand with OD pair $i$ and $j$ at time $t$
$G_{i}(n_{i}(t))$ Sum of trip completion and transfer flow in region $i$ at time $t$
$M_{ij}(t)$ Trip completion with OD from region $i$ to $j$ $(i\neq j)$ at time $t$
$n_{i,\rm cap}(t)$ Maximal number of vehicles (jam accumulation) in region $i$ at time $t$
$n_{i,\rm crit}(t)$ Vehicle accumulation with highest completion rate in region $i$ at time $t$
$J$ Objective function
2. Notations in reinforcement learning
$\mathcal{S}$ State space, the whole set of states the RL agent can transition to
$s_{t}$ $s_{t}\in\mathcal{S}$, the observable state in simulation at time $t$
$\mathcal{A}$ Action space, the whole set of actions the RL agent can act out
$a_{t}$ $a_{t}\in\mathcal{A}$, the action taken in simulation at time $t$
$\mathcal{R}$ The reward function for the RL agent
$r_{t}$ $r_{t}=\mathcal{R}(s_{t},a_{t})$, the received reward with state $s_{t}$ and action $a_{t}$ at time $t$
$\gamma$ Discount factor to favor rewards in the near future
$Q(s_{t},a_{t})$ Expected long-term return for taking action $a_{t}$ in state $s_{t}$ at time $t$
$\theta^{\mu}$ Weight parameter of the deep neural network for the actor network
$\theta^{Q}$ Weight parameter of the deep neural network for the critic network
$y_{i}$ Expected long-term return calculated with the target critic network
$L$ The loss of the critic network
$\rho^{\beta}$ All possible trajectories of $s_{t}$
$I$ The objective function for the actor-network
3. Notations for the additional antifragile terms applied in reinforcement learning
$r_{com}(t)$ Reward term in RL based on trip completion, equaling to $\sum^{2}_{i=1}M_{ii}(t)$
$r_{red}(t)$ Additional reward term in RL based on derivatives and redundancy
$r_{dam}(t)$ Additional damping term to counter possible action oscillation
$\omega_{h}$ The weight of first derivative in the additional reward term $r_{red}(t)$
$\omega_{\Delta h}$ The weight of second derivative in the additional reward term $r_{red}(t)$
$\alpha_{i}(t)$ Binary variable determining the term to be reward/penalty
$h_{i}(t)$ The first derivative of traffic state at time $t$
$\Delta h_{i}(t)$ The second derivative of traffic state at time $t$
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region, denoted as $q_{ii}(t)$, together with the perimeter control regulated transfer flow from region $j$ to region $i$, defined as $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by the trip completion within region $i$, denoted as $M_{ii}(t)$. similarly, the change rate of interregional traffic accumulation, as in Eq. 1b shows, is the difference between the interregional traffic demand, denoted as $q_{ij}(t)$, and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
$\frac{dn_{ii}(t)}{dt}$ $=q_{ii}(t)+u_{ji}(t)\cdot M_{ji}(t)-M_{ii}(t)$ (1a)
$\frac{dn_{ij}(t)}{dt}$ $=q_{ij}(t)-u_{ij}(t)\cdot M_{ij}(t),\>(i\neq j)$ (1b)
The total trip completion, i.e., $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD, defined as $G_{i}(n_{i}(t))$, and is the sum of the intraregional trip completion, i.e., $M_{ii}(t)$, in Eq. 2a and the interregional transfer flow, i.e., $M_{ij}(t)\>(i\neq j)$, in Eq. 2b:
$M_{ii}(t)$ $=\frac{n_{ii}(t)}{n_{i}(t)}\cdot G_{i}(n_{i}(t))$ (2a)
$M_{ij}(t)$ $=\frac{n_{ij}(t)}{n_{i}(t)}\cdot G_{i}(n_{i}(t)),\>(i\neq j)$ (2b)
$n_{i}(t)$ $=\sum_{j=1,2}n_{ij}(t)$ (2c)
The objective function is to maximize the throughput of this cordon-shaped network, which is the sum of the intraregional trip completion in both regions.
$J=\max\limits_{u_{ij}(t)}{\int_{0}^{T}\sum_{i=1,2}M_{ii}(t)dt}$ (3)
s.t. $n_{ij}(t)\geq 0$ (3a)
$n_{i}(t)\leq n_{i,\rm cap}$ (3b)
$u_{\min}\leq u_{ij}(t)\leq u_{\max}$ (3c)
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in the region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is necessitated since an indefinite long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the objective function for the proposed antifragile RL-based algorithms $J_{anti}$ is:
[…]
with $r_{com}(t)$ standing for the same trip completion as in Eq. 3:
$r_{com}(t)=\sum_{i=1,2}M_{ii}(t)$
The second term $r_{dam}(t)$ is the damping term to counter the oscillating behavior of the RL agents due to the incorporation of derivatives in RL. The third term $r_{red}(t)$ represents the redundancy term that aims to build up a proper redundancy so that the proposed RL algorithm does not reward the agent for targeting the optimal critical accumulation. More explanation of the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$ can be found in Section 4.2.
4 Methodology
In Section 2, we reviewed prior studies related to traffic control on inducing robustness and resilience using different RL algorithms. Here, we demonstrate how we integrate the concept of antifragility and enhance the performance of traffic control algorithms.
4.1 RL algorithm
In RL algorithms, an agent or multiple agents interact with a preset environment and improve the performance of decision-making, defined as action $a_{t}$ in an action space $\mathcal{A}$, based on the observable state $s_{t}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{t}=\mathcal{R}(s_{t},a_{t})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as the function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 2.
Figure 2: DDPG scheme
The actor-network is represented by $\mu(\cdot)$, and it determines the best action $a_{t}$ for the perimeter controller based on the current state $s_{t}$. The critic network $Q(\cdot)$ is responsible for evaluating whether a specific state-action pair at time step $t$ yields the maximal possible discounted future reward $Q(s_{t},a_{t})$. A common technique used in DDPG is to create a target actor network $\mu^{\prime}(\cdot)$ and a target critic network $Q^{\prime}(\cdot)$, which are a copy of the original actor and critic network but updated posteriorly to stabilize the training process and prevent overfitting (Zhang et al., 2021), with the target maximal discounted future reward calculated as:
$y_{i}=r_{i}+\gamma Q^{\prime}(s_{i+1},\mu^{\prime}(s_{i+1}|\theta^{\mu^{\prime}})|\theta^{Q^{\prime}})$ (5)
The critic network can then be updated by calculating the temporal difference between the predicted reward and the target reward and minimizing the loss for a mini-batch $N$ sampled from the replay buffer:
$L=\frac{1}{N}\sum_{i}(y_{i}-Q(s_{i},a_{i}|\theta^{Q}))^{2}$ (6)
Afterward, the actor network can be updated with sampled deterministic policy gradient:
$I=$ $\mathbb{E}_{s_{t}\sim\rho^{\beta}}[r(s,\mu(s|\theta^{\mu}))|_{s=s_{t}}]$ (7a)
$\nabla_{\theta^{\mu}}I=$ $\mathbb{E}_{s_{t}\sim\rho^{\beta}}[\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{t},a=\mu(s_{t})}\nabla_{\theta_{\mu}}\mu(s|\theta^{\mu})|_{s=s_{t}}]$ (7b)
$\nabla_{\theta^{\mu}}I\approx$ $\frac{1}{N}\sum_{t}\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{t},a=\mu(s_{t})}\nabla_{\theta^{\mu}}\mu(s|\theta^{\mu})|_{s_{t}}$ (7c)
First developed in Horgan et al. (2018), a similar Ape-X architecture as applied in Zhou and Gayah (2021) is also adopted in this work, which allows for multiple generations of simulations in the training process to be included and centrally learned for the best policy. A decaying Gaussian noise is also introduced to encourage the DDPG agent to explore the state space.
When training the agent, we add disruptions into the simulation environment starting from a certain training episode, such as surging traffic demand or MFD disruption. These various forms of volatility ought to optimally elicit various learning and decision-making processes. The RL agent would include the results of their prior choices to create and update their weight parameters in a situation with high outcome volatility and be capable of generating and updating expectations after sensing a change in a high-volatility environment.
4.2 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, robustness and resilience can be induced by integrating proper information into the RL algorithms. Aligning with the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the state $s_{t}\in\mathcal{S}$ is defined distinctively in different methods applied in this work. The benchmark work in Zhou and Gayah (2021) establishes the state as $s_{t}=[n_{ij}(t),q_{ij}(t)]$, with $q_{ij}(t)$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij}(t)$ can hardly be acquirable, particularly under an environment with disruptions. Therefore, we propose a state representation incorporating the derivatives of vehicle accumulation, which can be computed in the discrete form as:
$s_{t}=[n_{ij}(t),\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ (8)
The action $a_{t}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}(t)$. For the reward $r_{t}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$, as Eq. 4 shows. When $\Delta n_{ij}(t)$ and $\Delta^{2}n_{ij}(t)$ are incorporated into the state $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have little impact on certain performance indicators such as trip completion, the operation of traffic lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam}(t)$ is introduced into the reward function to penalize potential oscillatory actions. The exponent $\xi$ determines how fast the penalty decays when the change of control variables becomes small, indicating a large $\xi$ only penalizes the agent when the oscillation is substantial.
$r_{dam}(t)=-\sum_{i,j\in\{1,2\}}|u_{i,j}(t)-u_{i,j}(t-1)|^{\xi},\quad i\neq j$ (9)
The last term $r_{red}(t)$ in the objective function $J_{anti}$ acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling. However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivates similar to $\Delta n_{ij}(t)$ and $\Delta^{2}n_{ij}(t)$ in the state $\mathcal{S}$. Here, we summarize $r_{red}(t)$ as the sum of two terms $r_{red}(t)=H(t)+\Delta H(t)$, with $H(t)$ being an overall term representing the first derivative and $\Delta H(t)$ representing the second derivative, which can be further expanded as:
$H(t)=$ $\sum_{i=1,2}H_{i}(t)=\sum_{i=1,2}\omega_{h}\cdot h_{i}(t)\cdot\alpha_{i}(t)\cdot f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ (10a)
$\Delta H(t)=$ $\sum_{i=1,2}\Delta H_{i}(t)=\sum_{i=1,2}\omega_{\Delta h}\cdot\Delta h_{i}(t)\cdot f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ (10b)
Since normalization is common practice in RL to avoid gradient explosion, $\omega_{h}$ and $\omega_{\Delta h}$ are introduced as the weight constants for the derivatives to regulate their impact on the reward $\mathcal{R}$. $h_{i}(t)$ and $\Delta h_{i}(t)$ are the first and second difference of the traffic states on the MFD, $h_{i}(t)$ is defined as the difference of trip completion over vehicle accumulation at the end of a time step versus at the beginning of the same time step, and the second derivative $\Delta h_{i}(t)$ is calculated as the difference between the first differences of two consecutive time steps:
$h_{i}(t)=\frac{M_{i}(t)\ -\ M_{i}(t-1)}{n_{i}(t)\ -\ n_{i}(t-1)}$ (11a)
$\Delta h_{i}\left(t\right)=h_{i}\left(t\right)-h_{i}\left(t-1\right)$ (11b)
The binary variable $\alpha_{i}(t)$ in $H(t)$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent accumulation decreases.
$\alpha_{i}(t)=\begin{cases}1,&\text{if }n_{i}(t)\geq n_{i}(t-1),\\ -1,&\text{otherwise.}\end{cases}$ (12)
$f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red}(t)$ term when the accumulation around the critical accumulation $n_{i,\rm crit}$, where the $r_{red}(t)$ term should have the greatest impact. We use a modified trigonometric function to realize this purpose. Other functions, such as normal distribution, are also valid for achieving the same purpose.
$f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})=\begin{cases}{\Big{(}1+\cos\Big{(}-\pi\cdot\cfrac{n_{i,\rm crit}-n_{i}(t)}{n_{i,\rm crit}}\Big{)}\Big{)}}/2,&\text{if }n_{i}(t)\geq n_{i,\rm crit},\\ {\Big{(}1+\cos\Big{(}-\pi\cdot\cfrac{n_{i}(t)-n_{i,\rm crit}}{n_{i,\rm cap}-n_{i,\rm crit}}\Big{)}\Big{)}/2},&\text{otherwise.}\end{cases}$ (13)
We illustrate the impact of $H(t)$ and $\Delta H(t)$ on building redundancy on the MFD in Fig. 3 and 4. The first derivative $H(t)$ in Fig. 3(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n$ exceeds $n_{i,\rm crit}$. In Fig. 3, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.
An interesting note is that estimation uncertainty, also known as second-order uncertainty, is another factor that affects disruptions that take place unexpectedly. This is the imprecision of the learner’s current beliefs about the environment and what the antifragile terms capture. Estimation uncertainty reduces with sampling if beliefs are acquired by learning as opposed to instruction, such as anticipation through redundant overcompensation.
(a) $H(t)$ for the first difference $h_{i}(t)$.
(b) $\Delta H(t)$ for the second difference $\Delta h_{i}(t)$.
Figure 3: Illustration of the term $H(t)$ and its effect on the MFD
The second derivative $\Delta H(t)$ is shown in Fig. 4. The y-axis represents how fast the traffic state changes, and the faster it reaches $n_{i,\rm crit}$, the greater the penalty will be. This observation is consistent with the redundant overcompensation and time-scale separation principles formalized in Taleb and Douady (2013) and practically applied in Axenie and Saveriano (2023). On the contrary, a reward will be given if $n$ decelerates when approaching $n_{i,\rm crit}$. Likewise, $\Delta H(t)$ is also dependent on the normalization factor $\omega_{\Delta h}$. With $H(t)$ and $\Delta H(t)$, the agent learns to be conservative when regulating the perimeter control variables to reach $n_{i,\rm crit}$.
Figure 4: Illustration of the term $\Delta H(t)$
4.3 Model Predictive Control
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in Model Predictive Control (MPC) and its applications in traffic engineering can refer to Geroliminis et al. (2013); Haddad and Mirkin (2017). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019) and the NLP solver IPOPT (Wächter and Biegler, 2006).
5 Experiment application
This section introduces the most important settings for the numerical simulation to showcase antifragile perimeter control and the corresponding performance evaluation. In addition to No Control (NC), we investigated two state-of-the-art perimeter control algorithms compared to our proposed antifragile algorithm, which are summarized below:
    NC where the system operates without intervention.
    MPC proposed in Geroliminis et al. (2013).
    State-of-the-art RL-based perimeter control algorithm proposed in Zhou and Gayah (2021) as the RL benchmark.
The performance of the traffic control algorithms is studied under two different types of disruptions as illustrated in Fig. 5, i.e., a demand disruption like surging traffic demand, or a supply disruption, such as the change of MFD for a specific network due to adversarial weather or major accidents. Furthermore, to showcase both proto-antifragility and progressive antifragility, this work examines the impact of different manifestations of disruption magnitudes. In the short term, the magnitudes are assumed to remain relatively stable with minor stochastic variations. In the long term, however, as traffic demand continues to increase and contributes to more frequent traffic incidents, the magnitude is also expected to grow, which can be subject to real-world uncertainties as well. The static and incremental magnitude of disruptions is intended for testing proto- and progressive antifragility. To sum up, the combination of 1) demand and supply disruptions, 2) static and incremental disruptions, and 3) with and without uncertainty in disruptions has resulted in a total number of 8 scenarios.
Figure 5: Illustration of simulation scenarios and their evaluation
5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 1(b) shows. These MFDs are largely the same as in Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, and indifferentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by $3\%$, so that the whole MFD can be both continuous and differentiable. Other important indicators, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
Even though real-world disruptions can happen to both the outer and the inner regions and exert a negative influence, we focus on the bottleneck scenarios where the potential of perimeter control can be better reflected. Therefore, for a demand disruption, a surging traffic demand is assumed to be generated from the outer region toward the inner region, whereas for a supply disruption, the inner MFD is lowered as a result of disruptive events while the outer MFD remains intact. Demand and supply disruptions also demonstrate different training difficulties for the two RL-based algorithms involved, which will be detailed in Section 6.
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 6(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.3,0.35,0.2,0.25]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,N}(t)=C_{ij,N}\mathcal{N}_{ij}(\mu_{ij,N},\sigma_{ij,N})$ with total number of vehicles following Gaussian as $C_{ij,N}=[3000,15000,2000,4000]$ and the mean and variance of the distribution being $\mu_{ij,N}=[1800,1800,1800,1800]$ and $\sigma_{ij,N}=[1200,1500,900,1200]$ in seconds. Under a demand disruption scenario, $q_{12,N}$ is increased to simulate traffic flowing from the periphery into the city center. For testing progressive antifragility, a compounding factor of $0.018$ per episode is applied to escalate $q_{12,N}$, with the maximal demand disruption after 25 episodes reaching $(1.018^{25}-1)C_{12,N}\mathcal{N}_{12}(\mu_{12,N},\sigma_{12,N})$. On the other hand, the magnitude for validating proto-antifragility is set to be static as $(1.018^{25/2}-1)C_{12,N}\mathcal{N}_{12}(\mu_{12,N},\sigma_{12,N})$.
(a) Base demand profile and disruptions
(b) MFD and disruptions for the inner region
Figure 6: Demand profile with or without surging demand
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. Recent work in Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is used considering the performance loss due to vehicle infrastructure interaction, and a smaller $\lambda$ indicates the infrastructure is utilized with higher efficiency. In this research, a $\lambda$ equal to $0.07$ is chosen for the MFD without disruption representing the inner region, which lies in the common range of $\lambda$ based on data from real-world cities (Ambühl et al., 2020). And for the incremental magnitude of supply disruptions to test progressive antifragility, $\lambda$ is escalated by $0.003$ per episode with the maximal $\lambda$ after 25 episodes being $0.145$, as the solid red curve shows in Fig. 6(b). For static supply disruptions to test proto-antifragility, the magnitude of disruption is set to be half of the maximal disruption. Note that since the polynomial MFD in Geroliminis and Daganzo (2008) and the MFD approximated with $\lambda$ are essentially two different functional forms, there are still marginal deviations between each other after careful parameter calibration, as illustrated with the blue and gray curves in Fig. 6(b).
The total simulation duration $T$ is 3 hours with each time step $\Delta t$ as $180$ seconds, and the third hour has little traffic demand and mainly serves as the unloading process. Another notable constraint is the lower and upper bounds for the perimeter control variable $u_{ij}(t)\in[0.1,0.9]$ to represent the real-world constraints from traffic signals (Geroliminis et al., 2013). The initial vehicle accumulation is set to be $n_{ij}(0)=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation.
Each scenario is run $25$ times since randomness is inherent in RL, and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Either static or incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 5. A list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The multiplier list will be shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through reshuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers but with different sequences.
The most important hyperparameters for both the baseline and the proposed RL algorithms should be the same and are summarized in Table. 2 below. Note that the minimal learning rates and noise scale are not set to be an extremely small value as common RL algorithms do, we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions, which can be demonstrated by the results for both good performance and antifragility in Section 6. The exponent coefficient $\xi$ for the damping term $r_{dam}(t)$ in Eq. 9 to counter the potential oscillations actions is set to be $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red}(t)$ is set to be $0.1$ and $0.2$. The number of simulations generated under the Ape-X architecture is 32 per training episode.
Table 2: List of hyperparameters
Hyperparameter Value
Replay buffer 10,000
Sample size 1,000
Action noise initial scale 0.3
Action noise linear decay 0.003
Action noise minimal scale 0.1
Batch size 256
Target network update 5
Discount factor 0.90
Actor Critic
Initial learning rate 0.004 0.008
Learning rate decay 0.975 0.975
Minimal learning rate 0.001 0.002
Epoch 2 128
5.2 Performance evaluation
The reward in the RL algorithm is defined based on trip completion with additional damping redundancy terms. However, two scenarios with the same trip completion at the end of the simulation may exhibit distinct TTS, and the one with the lower value should be regarded to have demonstrated a superior overall performance. In addition, to better showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 13, the main performance indicator being evaluated in this work is the Total Time Spent (TTS), which is calculated by adding up the number of vehicles within the network at each second of the simulation.
Since urban road networks are always subject to capacity constraints, and as Sun et al. (2024) has recently proven the fragile nature of urban road networks in terms of progressive fragility, fully antifragile traffic control strategies may be impossible to achieve and will inevitably fall into being fragile when the magnitude of disruptions is large enough. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is less fragile than the state-of-the-art baseline algorithms, and we normalize all the other perimeter control algorithms over the RL baseline method to study the relative antifragility. To quantify the progressive antifragility of different algorithms, we calculate the skewness of distribution based on the samples from the last 25 incremental episodes, with $\mu$ and $\sigma$ denoting the mean and the standard deviation:
$s=\frac{1}{N_{episode}}\sum_{i=1}^{N_{episode}}\left(\frac{TTS_{i}-\mu}{\sigma}\right)^{3}$ (14)
A skewness computed to be 00 means that the system itself or the applied algorithm makes the system neither fragile nor antifragile. On the other side, negative skewness indicates the distribution has a longer or fatter right tail and thus a higher degree of concavity in the performance function, which showcases antifragility, whereas a positive skewness signifies fragility. Therefore, other than demonstrating a superior performance regarding a lower TTS, the proposed method should also showcase a lower skewness than the baseline method.
6 Results
As mentioned in the previous section, there are 8 simulation scenarios based on the combination of demand/supply disruption, static/incremental disruption, and with/without uncertainty, with each one containing the performance of 4 perimeter control algorithms to be tested: NC, MPC, the baseline RL algorithm (Zhou and Gayah, 2021), and our proposed RL-based algorithm. Here, we first present the results of simulations with a static magnitude of disruptions to showcase proto-antifragility, followed by the results with incremental magnitudes to demonstrate progressive antifragility. Each scenario is tested for 25 simulation iterations.
6.1 Proto-antifragility
Systems with the property of proto-antifragility can learn from past adverse events and anticipate possible ongoing disruptions of similar magnitudes to enhance future performance. To validate such property for the RL-based algorithms, after training the agent with the base demand profile for 50 episodes, we apply disruptions with static magnitude for another 25 episodes.
Fig. 7 demonstrates the performance curves of the four algorithms under no uncertainty or with uncertainty. Each data point on any performance curve for a particular episode represents the TTS averaged from 25 simulation iterations of 3 hours. The shaded area represents the performance variance due to the innate randomness in the training process of the RL agents or combined with uncertainties introduced following the shuffling scheme described in Section. 5.1.
(a) Without disruption uncertainty
(b) With disruption uncertainty
Figure 7: Performance curves under static demand disruptions
In the 50 episodes of training with the base demand profile in Fig. 7(a), the proposed RL algorithm outperforms both NC and the benchmark RL algorithm and demonstrates performance comparable to that of MPC. In the following 25 episodes with disruption, both RL-based algorithms exhibited improved performance following an initial rise of TTS at the transition and outperform NC and MPC. Overall, the proposed RL algorithm performs better than the baseline RL algorithm before the introduction of the static disruption and shows an even superior performance of about $4.2\%$ after the disruption comes into play. In addition, the performance variance of the proposed RL algorithm is significantly smaller than the benchmark RL algorithm. In Fig. 7(b), as disruption magnitude is set to be stochastic following a Gaussian distribution, the performance of NC and MPC also contains a shaded area. The shaded area is rather stable under uncertainty due to the reshuffling mechanism explained in Section 5 to reshuffle the uncertainty multiplier list $[\epsilon_{i}]$ by $1$ after each iteration. Despite the enlarged variance for both RL-based methods, the averaged TTS curves are largely the same as in 7(a) with no uncertainty. The performance gain of the proposed RL algorithm compared to the RL baseline under disruption uncertainty is $3.9\%$, marginally smaller compared to the performance without uncertainty.
Likewise, as demonstrated in Fig. 8, similar results can be observed in the study of static supply disruption. The averaged TTS curves show a $3.0\%$ performance gain of the proposed RL algorithm compared to the baseline RL algorithm under no disruption uncertainty. In Fig. 8(b), the performance gain is marginally higher and comes to $3.5\%$ under the presence of disruption uncertainty. However, the performance variance of the proposed algorithm in Fig. 8(a) is greater with disruption uncertainty compared to the baseline RL algorithm, but still, its shaded area in blue lies predominantly under the gray shaded area, indicating the superior performance of the antifragile RL algorithm.
(a) Without disruption uncertainty
(b) With disruption uncertainty
Figure 8: Performance curves under static supply disruptions
6.2 Progressive antifragility
Progressive antifragility describes a system that exhibits a nonlinear response to growing disruptions. Since an increase in traffic density or vehicle accumulation within a network will inevitably reduce average vehicle speed or trip completion and increase TTS, a concave response between TTS and vehicle accumulation can indicate a progressively antifragile system. Instead of 25 episodes of a static disruption, a linearly increasing magnitude of demand or supply disruption is simulated, and the distribution skewness is calculated to determine system (anti-)fragility.
6.2.1 Demand disruption
In Fig. 9(a), when no disruption uncertainty is present in the last 25 episodes of incremental disruptions, the proposed RL algorithm performs the best out of all four tested algorithms. The shaded area represents the variance from the 25 iterations of the simulation and is lower than the RL baseline. Furthermore, the performance curve of the proposed method also seems less convex than the other algorithms, demonstrating its relative antifragility. Before demonstrating antifragility quantitatively, it should be noted that the TTS curves from NC and MPC are only partially illustrated, as their simulated network came into gridlock quicker when faced with a major disruption with the vehicle accumulation reaching maximal accumulation. And when the network is fully gridlocked, it cannot recover to an uncongested state even when the surging demand has declined. In this case, TTS will only grow linearly with the magnitude of distribution and lose the ability to showcase skewness, which will be further detailed below in Fig. 9(c).
(a) Performance curves in TTS
(b) Normalized over Zhou & Gayah (2021)
(c) Skewness up to the $n$-th episode
Figure 9: Performance curves under incremental demand disruptions
(a) Performance curves in TTS
(b) Normalized over Zhou & Gayah (2021)
(c) Skewness up to the $n$-th episode
Figure 10: Performance curves under incremental demand disruptions with uncertainty
In Fig. 9(b), the performance difference normalized over the baseline RL algorithm is illustrated with a polar plot. Only the episodes $50-75$ under incremental demand are shown. Instead of TTS variance, the shaded blue area highlights the performance gain of our proposed method compared to the state-of-the-art baseline RL algorithm. As observed, the TTS curves for NC, MPC, and the proposed RL algorithm all diverge from the baseline method at an accelerating rate, indicating diverse degrees of relative (anti-)fragility. The main comparison here is the two RL-based algorithms, and our proposed algorithm shows an ultimate $9.3\%$ performance gain after 25 episodes with incremental demand disruption. Also, the disruption magnitude between the 62-rd and 63-rd episodes under incremental disruption is about the same as in the simulation with static disruption, and it is interesting to see that the performance gain here is approximately $5.2\%$, significantly higher than the $4.2\%$ improvement with a static magnitude of disruption, hinting the proposed method adapts relatively faster under an incremental environment.
To analyze the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. 9(c), so that the skewness value at the final episode reflects the distribution skewness across all $25$ episodes with incremental disruptions. Since calculating skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase. Initially, the skewness for all four methods starts at a low value. However, the skewness for NC and MPC increases rapidly to approximately $1$, whereas the skewness for the two RL-based algorithms rises more slowly. Among these, the skewness curve in blue of the proposed algorithm lies mostly below the baseline curve in gray, particularly when the magnitude of disruption is high, achieving the lowest final skewness at $0.29$. It should be noted that the results from both NC and MPC indicate that these two algorithms render the network quickly into gridlock under the presence of demand disruptions, where the TTS will only grow linearly with the magnitude of disruptions and yield a skewness of 00. Therefore, we use dashed lines to represent the maximal skewness if the skewness of a particular algorithm has ever plateaued.
When uncertainty is introduced to the magnitude of disruptions, as shown in Fig. 10, the results remain largely consistent with the non-stochastic scenario, with the performance curves being less smooth, suggesting that the presence of uncertainty adds complexity to the training process. Compared to the baseline algorithm, the proposed RL algorithm demonstrates a continuously increasing performance gain, as shown in Fig. 10(b), reaching $9.2\%$ when the magnitude of disruption becomes significant. The unevenness in performance is also reflected in the value of skewness in Fig. 10(c), despite the final skewness of the two RL-based algorithms $0.32$ and $0.29$ being almost the same $0.32$ and $0.28$ as in Fig. 9(c). The final skewness values for NC and MPC are noticeably lower. The introduction of uncertainty causes gridlock to occur sooner, resulting in the premature entering into plateauing for both the NC and MPC, particularly when considering the convex fragile performance of the network.
6.2.2 Supply disruption
TTS curves of the four studied algorithms under deterministic supply disruption are shown in Fig. 12. The magnitude of incremental supply disruption is tuned so that the performance curve from NC is largely the same as with incremental demand disruption. However, the TTS from MPC and RL-based algorithms has somewhat increased, showing an increase of roughly $15\%$. The difference in performance has demonstrated the distinct levels of difficulty demand and supply disruption have posed for the learning of the RL agents. The relatively faster-growing skewness in Fig. 11(c) is yet another evidence of such challenges. The state of the baseline RL algorithm is composed of $[n_{ij}(t),q_{ij}(t)]$, whereas it is $[n_{ij}(t),\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ in the proposed algorithm. Although the baseline method relies on additional information of $[q_{ij}(t)]$, it becomes misinformation under demand disruption, whereas the derivatives $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ have a similar function as $[q_{ij}(t)]$ but being less susceptible to demand disruption. When it comes to disruption on the supply side, the situation differs because neither of the state definitions of the two RL-based algorithms is capable of accurately reflecting the model shifting. Still, with a faster changing rate of the derivatives, the agents in the proposed algorithm can make better decisions and gradually adapt to the new model. As shown in Fig. 11(b), the performance gain can grow up to $12.2\%$ when the supply disruption is substantial with $\lambda$ being $0.145$, demonstrating the superior performance of the proposed method compared to the baseline method. Furthermore, in Fig. 11(c), the blue skewness curve lies constantly below the baseline curve in gray, with an ultimate skewness of $0.41$, which indicates that the proposed RL algorithm is relatively more antifragile than the baseline algorithm.
(a) Performance curves in TTS
(b) Normalized over the baseline RL algorithm
(c) Skewness up to the $n$-th episode
Figure 11: Performance curves under incremental supply disruptions
(a) Performance curves in TTS
(b) Normalized over the baseline RL algorithm
(c) Skewness up to the $n$-th episode
Figure 12: Performance curves under incremental supply disruptions with uncertainty
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. 12, the results are similar to those observed under demand disruption with uncertainty. First, the TTS difference is $11.8\%$, showing minimal variation between scenarios with and without disruption. The skewness of the two RL-based algorithms remains almost unchanged, shifting only slightly from $0.41$ and $0.52$ to $0.42$ and $0.51$ for the proposed and baseline algorithms, respectively. In contrast, the skewness values for NC and MPC decrease from $0.88$ and $0.78$ to $0.82$ and $0.73$, reflecting the fact that the traffic state has already reached the gridlock accumulation without uncertainty, and by the incorporation of uncertainty, the network is prone to get gridlocked even more easily. Likewise, even with an average of 25 simulation iterations, the performance curves under supply disruption with uncertainty are less smooth, which is a similar pattern under demand disruption uncertainty.
7 Conclusion
This work introduces the concept of antifragility by comparing it with two other terms, i.e., robustness and resilience, which are commonly used in transportation and traffic control strategies. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust or resilient designs, we explore whether and how these properties can be induced with RL. Building on this concept, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives — representing the rate of change and curvature of traffic states — to provide the RL algorithm with richer information. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions.
We conducted comprehensive experiments to compare our proposed antifragile RL-based perimeter control approach against three other methods: NC, MPC, and the state-of-the-art RL-based method as the baseline. This evaluation was performed on two levels—disruptions with static and incremental magnitudes to validate proto-antifragility and progressive antifragility, respectively. In addition, we examined two distinct manifestations of disruptions: demand disruptions and supply disruptions. Uncertainties in the magnitude of disruptions are also considered. The results demonstrated the effectiveness of our proposed method based on superior performance regarding lower TTS. More importantly. The proposed algorithm showcased different levels of antifragility under static and incremental disruptions. We put forward a novel method for quantifying antifragility by computing distribution skewness. The proposed RL-based antifragile algorithm exhibits the lowest skewness among all the methods examined, indicating its progressive antifragility property compared to the other examined methods.
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation of transportation networks, improving performance under unforeseen disruptions with learning-based algorithms. In addition, it introduces a new possibility for evaluating system operation under disruptive conditions. Moreover, the concept is generic to be extended to other traffic control systems and potentially to systems that are subject to disruptions in general.
8 CRediT authorship contribution statement
Linghang Sun: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft. Michail A. Makridis: Conceptualization, Methodology, Supervision, Writing - review & editing. Alexander Genser: Methodology, Visualization, Writing - review & editing. Cristian Axenie: Project administration, Resources, Writing - review & editing. Margherita Grossi: Project administration, Resources, Writing - review & editing. Anastasios Kouvelas: Supervision, Writing - review & editing.
9 Acknowledgement
This research was kindly funded by the Huawei Munich Research Center under the framework of the Antigones project.
10 Declaration of Competing Interest
This research was kindly funded by the Huawei Munich Research Center under the framework of the Antigones project, with one of our co-authors being employed at the said company. Otherwise, the authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Graphical illustration of robustness, resilience, and antifragility
Fig. 13 graphically illustrates the differences in robustness, resilience, proto-antifragility, and progressive antifragility. In brief, robustness is about the system’s resistance to minor disturbances, whereas resilience represents the ability to recover from major disruptions. Proto-antifragility enables the system to adapt to disruptions with magnitudes falling within a similar range, while the performance of a progressive antifragile system exhibits a concave response with linearly growing disruptions, either due to the inherent properties of the system or through antifragile control strategies.
Figure 13: Illustration for robustness, resilience, adaptiveness, and antifragility