### v5

# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies reliant on mathematical models can struggle with real-world disruptions, leading to significant divergence from their anticipated efficiency. This work fuses the cutting-edge concept of antifragility into the design of a learning-based traffic control strategy for large-scale urban road networks to perform against disruptions. Antifragility sets itself apart from robustness, resilience, and reliability as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile modules composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network, and a numerical case study based on real-world data. Promising results highlight (1) the superior performance of the proposed algorithm compared to the state-of-the-art methods under incremental magnitude of disruptions, (2) distribution skewness as the antifragility indicator demonstrating its relative antifragility, (3) its effectiveness under limited observability due to real-world data availability constraints, and (4) robustness of the proposed antifragile module when embedded into advanced RL algorithms for superior performance. The proposed antifragile methodology is generalizable and holds potential for applications beyond perimeter control, offering integration into systems exposed to disruptions across various disciplines.
keywords:
antifragility , reinforcement learning , perimeter control , demand and supply disruption , macroscopic fundamental diagram
journal: Reliability Engineering & System Safety
[…]
1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations. Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness, resilience, reliability, and even safety of transportation systems is highly crucial (Ganin et al., 2019; Zayandehroodi et al., 2025). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Albalate and Fageda, 2021). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with linearly increasing magnitudes of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and the exponentially growing negative effects.
To address such issues, the concept of antifragility has shed light on a potential solution. Echoing the sentiment of the well-known maxim What does not kill me makes me stronger in Nietzsche (1889), antifragility was first introduced in the reputable general-audience publication Antifragile: Things That Gain from Disorder (Taleb, 2012). It has then been mathematically formulated as an academic concept in Taleb and Douady (2013), providing insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks, such as in transportation networks (Cats and Hijner, 2021). Ever since the concept was proposed, antifragility has gained substantial interest from both the public and academia, particularly in the complex systems and risk engineering communities (Aven and Krohn, 2014; Aven, 2015; Thekdi and Aven, 2019; Aven, 2022; Axenie et al., 2024). However, the design principles for realizing antifragile transportation systems and control are generally unexplored. One promising approach to induce antifragility is to leverage learning-based algorithms, such as Reinforcement Learning (RL), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
The primary goal of this paper is to design an antifragile perimeter control algorithm capable of adapting to increasing magnitudes of disruptions, reflecting the effects of urban densification and rising traffic demand. Fig. 1 presents a map of contributions while highlighting the key references. The contributions of this research can be summarized as follows:
  1. 1.
    Conceptual contribution: We distinguish the concept of antifragility as opposed to other related and commonly used terms in the risk engineering and transportation domains.
  2. 2.
    Methodological contribution: We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods, tackling both fragile performance and observability issues.
  3. 3.
    Evaluatory contribution: We adopt a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under increasing disruptions.
  4. 4.
    General applicability: We further tested the effectiveness of the proposed antifragile module on other state-of-the-art RL algorithms and validated its applicability.
Figure 1: Map of contribution
The remainder of this paper is structured as follows. In Section 2, we introduce relevant literature on multiple aspects covered in this work, while Section 3 mathematically formulates the cordon-shaped perimeter control simulation environment. Methodologies related to incorporating antifragility into RL algorithms are detailed in Section 4. Section 5 discusses the simulation setups and parametrization. Results are presented in Section 6, followed by concluding remarks and further discussions in 7.
2 Literature review
This section reviews the relevant literature on three topics intertwined in this work. First, a macroscopic traffic model and the control strategy applied in this work are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we review how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, related terminologies to antifragility are distinguished while introducing the design philosophy of an antifragile system.
2.1 Macroscopic fundamental diagram and perimeter control
Alleviating urban network congestion can be realized through various traffic control strategies. Since Geroliminis and Daganzo (2008) demonstrated the presence of the Macroscopic Fundamental Diagram (MFD) with empirical data, the mathematical relationship between flow-density-speed has paved the path for the development of control strategies at a macroscopic level, enabling more computationally feasible real-time traffic control strategies for large-scale networks, such as perimeter control (Keyvan-Ekbatani et al., 2012; Geroliminis et al., 2013; Kouvelas et al., 2017), congestion pricing (Genser and Kouvelas, 2022), route guidance (Fu et al., 2022).
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Sirmatel and Geroliminis (2020) introduced a Moving Horizon Estimation (MHE) scheme together with MPC to tackle further measurement noise and the observability issues. One major issue of the previous works is the assumption of MFD homogeneity, and a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for each sub-network, as in Ambühl et al. (2019), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers. Therefore, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free.
2.2 Leveraging RL to induce robustness, resilience, and reliability
Similar to traffic networks and system control, real-world systems of all kinds can be subject to disruptions. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged such features to enhance system performance under adversarial events. Here, we present a review of studies that apply RL in reliability engineering and traffic engineering domains, claiming enhancements in robustness, resilience, reliability, or related properties.
Most existing studies on data-driven maintenance scheduling algorithms for infrastructure systems or mechanical equipment prioritize costs or efficiency, often lacking an orientation to reliability or resilience. To bridge such gaps, Yuan et al. (2025) proposed an innovative reliability reward component based on the Gamma-Wiener process, enabling the RL agent to achieve a balance between dynamic reliability and maintenance cost of mechanical equipment. Likewise, Liang et al. (2026) proposed a specialized resilience metric, which encourages the RL agent to accelerate post-disaster recovery while explicitly considering system resilience, while Yin et al. (2025) couples the DistFlow model for feasible resilience assessment and Bayesian learning for the quantification of system uncertainties to promote resilient post-disruption recovery. Risk and uncertainty can also be directly modeled in the reward function as in Chiou (2024). On top of reward shaping, Xu et al. (2026) also introduced a phase-adaptive discount factor to enhance the resilience of combat system-of-systems.
In traffic networks, RL algorithms can also be engineered to strengthen the robustness or resilience of different control systems, such as traffic light control or perimeter control. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are provided with additional information related to disturbances or disruptions.
The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evoke unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately, since they might indicate a shift in the environment. In this case, exploration refers to the process of searching for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness, resilience, and reliability can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between antifragility and other closely related terms, which include robustness, resilience, reliability, stability, and adaptiveness. We follow the definitions proposed in Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. Reliability in transportation has a much broader range of meanings. Pennetti et al. (2020) proposed travel time reliability, which has a similar flavor to robustness in measuring deviation from normal operation, but focuses on the probability of such deviation exceeding a certain threshold. Cats et al. (2017) introduced a reliability assessment indicator to account for performance loss for the entire range of possible capacity reductions, whereas robustness pertains merely to a single disruption event. Stability can also bear multiple definitions based on the context, such as describing the closed-loop stability of system control as in Sirmatel and Geroliminis (2021) or MFd hysteresis on the traffic flow theory basis (Gayah and Daganzo, 2011). Adaptiveness describes the ability of systems to alter their traits to satisfy autonomy in different environments (Hooker, 2011), which aligns with the concept of proto-antifragility proposed in Taleb (2012). However, adaptiveness does not necessarily guarantee a consistent performance improvement under growing magnitudes of disruptions. Therefore, Taleb (2012) introduced the concept of antifragility, emphasizing the concave response of the system under increasing disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is illustrated and described in Fig. 12. It should be noted that while (anti-)fragility can be an innate property of a system, proper intervention and control strategies can lead the system to be more antifragile (Axenie et al., 2024). For instance, building on the mathematical proof by Sun et al. (2024) that urban road networks are naturally fragile, this paper seeks to develop antifragile solutions that mitigate such fragility through perimeter control.
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines in recent years, such as energy (Coppitters and Contino, 2023), electricity (Rachunok and Nateghi, 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), cyber-systems (Chatterjee and Thekdi, 2020), robotics (Axenie and Saveriano, 2023), and machine learning (Pravin et al., 2024). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives of system states to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Axenie et al., 2022). With this additional information, the system can potentially anticipate ongoing disruptions and become more proactive in facing drastic changes. Similar to its function in resilience (Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, leveraging the potential of antifragile system design for the operation of transportation networks is still an unexplored notion.
3 Problem formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 2(a). The total number of vehicles in region $i$ at time $t$ is denoted as $n_{i}(t)$, while the Origin-Destination (OD) from region $i$ to region $j$ can be further divided as $n_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference of accommodating vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 2(b). The total trip completion rate $M_{i}(t)$ for region $i$ at time $t$ can be determined through the corresponding MFD, which comprises both the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. While $q_{i}(t)$ denotes the new demand within region $i$ at time $t$, $q_{ij}(t)$ represents the additional demand based on the OD. To protect a certain region from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to get across the perimeter is regulated by the perimeter controllers, denoted as $u_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. All notations applied in this work, including those in the following sections, are summarized in Table 1.
(a) The cordon-shaped urban network
(b) MFDs for the inner and outer regions.
Figure 2: The network structure and the related MFDs.
Table 1: List of notations
Symbol Description
1. Notations for the problem formulation (continuous form)
$t$ Current time
$\Delta t$ Time step interval
$T$ Total simulation time
$n_{ij}(t)$ Vehicle accumulation with OD from region $i$ to $j$ at time $t$
$n_{i}(t)$ Vehicle accumulation in region $i$ at time $t$
$u_{ij}(t)$ Perimeter control variables regulating flow from region $i$ to $j$ at time $t$
$q_{ij}(t)$ Traffic demand with OD pair $i$ and $j$ at time $t$
$G_{i}(n_{i}(t))$ Sum of trip completion and transfer flow in region $i$ at time $t$
$M_{ij}(t)$ Trip completion with OD from region $i$ to $j$ at time $t$
$n_{i,\rm cap}(t)$ Maximal number of vehicles (jam accumulation) in region $i$ at time $t$
$n_{i,\rm crit}(t)$ Vehicle accumulation with highest completion rate in region $i$ at time $t$
$J$ Objective function
2. Notations in RL (discrete form)
$k$ Current time step, $k=\lfloor t/\Delta t\rfloor$
$K$ Total time steps, $K=T/\Delta t$
$\mathcal{S}$ State space, the whole set of states the RL agent can transition to
$s_{k}$ $s_{k}\in\mathcal{S}$, the observable state in simulation at timestep $k$
$\mathcal{A}$ Action space, the whole set of actions the RL agent can act out
$a_{k}$ $a_{k}\in\mathcal{A}$, the action taken in simulation at timestep $k$
$\mathcal{R}$ The reward function for the RL agent
$r_{k}$ $r_{k}=\mathcal{R}(s_{k},a_{k})$, the received reward with state $s_{k}$ and action $a_{k}$ at timestep $k$
$\gamma$ Discount factor to favor rewards in the near future
$Q(s_{k},a_{k})$ Expected long-term return for taking action $a_{k}$ in state $s_{k}$ at timestep $k$
$\theta^{\mu}$ Weight parameter of the deep neural network for the actor network
$\theta^{Q}$ Weight parameter of the deep neural network for the critic network
$y_{i}$ Expected long-term return calculated with the target critic network
$L$ The loss of the critic network
$\rho^{\beta}$ All possible trajectories of $s_{k}$
II The objective function for the actor-network
3. Notations for the antifragile terms in RL (discrete form)
$r_{com,k}$ Reward term in RL based on trip completion
$r_{red,k}$ Additional reward term in RL based on derivatives and redundancy
$r_{dam,k}$ Additional damping term to counter possible action oscillation
$\omega_{h}$ The weight of the first derivative in the additional reward term $r_{red,k}$
$\omega_{\Delta h}$ The weight of the second derivative in the additional reward term $r_{red,k}$
$\alpha_{i,k}$ Binary variable determining the term to be a reward/penalty
$h_{i,k}$ The first derivative of traffic state at timestep $k$
$\Delta h_{i,k}$ The second derivative of traffic state at timestep $k$
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region $q_{ii}(t)$ with the perimeter control regulated transfer flow $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by trip completion $M_{ii}(t)$. Likewise, the change rate of interregional traffic accumulation, as Eq. 1b shows, is the difference between the interregional traffic demand $q_{ij}(t)$ and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
$\frac{dn_{ii}(t)}{dt}$ $=q_{ii}(t)+u_{ji}(t)\cdot M_{ji}(t)-M_{ii}(t)$ (1a)
$\frac{dn_{ij}(t)}{dt}$ $=q_{ij}(t)-u_{ij}(t)\cdot M_{ij}(t)$ (1b)
The total trip completion $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD $G_{i}(n_{i}(t))$, which is the sum of the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)$:
$M_{ii}(t)$ $=\frac{n_{ii}(t)}{n_{i}(t)}\cdot G_{i}(n_{i}(t))$ (2a)
$M_{ij}(t)$ $=\frac{n_{ij}(t)}{n_{i}(t)}\cdot G_{i}(n_{i}(t))$ (2b)
$n_{i}(t)$ $=\sum_{j=1,2}n_{ij}(t)$ (2c)
The objective function is to maximize the throughput of this cordon-shaped network, which is the sum of the intraregional trip completion $M_{ii}(t)$ in both the inner and the outer regions, while the interregional trip completion represents only vehicles finishing part of their trips.
$J=\max\limits_{u_{ij}(t)}{\int_{0}^{T}\sum_{i=1,2}M_{ii}(t)dt}$ (3)
s.t. $n_{ij}(t)\geq 0$ (3a)
$n_{i}(t)\leq n_{i,\rm cap}$ (3b)
$u_{\min}\leq u_{ij}(t)\leq u_{\max}$ (3c)
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ in Eq. 3c represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is essential since an indefinitely long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the reward function for the proposed antifragile RL-based algorithms $J_{af}$ is described in Eq. 4 in discrete form, with $r_{com,k}$ standing for the total trip completion in both regions at timestep $k$, that is $r_{com,k}=\sum\limits^{k\Delta t+\Delta t}_{t=k\Delta t}\sum\limits_{i=1,2}M_{ii}(t)$.
$J_{af}=\max\limits_{u_{ij}(k)}{\sum_{k=1}^{K}\left(r_{com,k}+r_{dam,k}+r_{red,k}\right)}$ (4)
The second term $r_{dam,k}$ is the damping term to counter the oscillating behavior of the RL agents due to the incorporation of derivatives in RL. The third term $r_{red,k}$ represents the redundancy term that aims to build up a proper redundancy so that the proposed RL algorithm does not reward the agent for targeting the optimal critical accumulation. More explanation of the damping term $r_{dam,k}$ and the redundancy term $r_{red,k}$ can be found in Section 4.3.
4 Methodology
In this section, we first discuss the methodologies of the benchmarks and use their benefits and drawbacks to motivate the development of our proposed antifragile perimeter control algorithm.
4.1 Model predictive control and moving horizon estimation
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in MPC and its applications in traffic engineering can refer to Geroliminis et al. (2013). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019).
In Sirmatel and Geroliminis (2020), Moving Horizon Estimation (MHE) is introduced to form an MPC-MHE framework for perimeter control. MHE enables real-time estimation of internal system states by solving an optimization problem that minimizes the discrepancy between predicted outputs and noisy sensor measurements. It further addresses the data observability issues in Geroliminis et al. (2013) as illustrated in Fig. 1, that is, the information on both the OD accumulation $n_{ij}(t)$ and the newly generated demand $q_{ij}(t)$ is almost impossible to acquire in real-time. Therefore, four different combinations of data availability have been tested, i.e., one from either $n_{ij}(t)$ or $n_{i}(t),M_{ij}(t)]$ and another one from either $q_{ij}(t)$ or $q_{i}(t)$. The accumulation $n_{i}(t)$ for region $i$ can be approximated with loop detector measurements, while interregional trip completion $M_{ij}(t)$ can be directly counted through detectors at the intersections forming the perimeter border. The authors also claim $q_{i}(t)$ can be available with some applications, such as a substantial percentage of GPS information being readily accessible. However, due to the nature of the prediction horizon in MPC-MHE, such data has to be made available in both real-time and the near future, rendering it impossible to collect, especially under disruptive conditions. We will further tackle such observability issues with the proposed antifragile perimeter control algorithm in the following Section 4.3.
4.2 The DDPG RL algorithm
In RL algorithms, an agent or multiple agents interact with a preset environment and improve its decision-making capacilities, defined as action $a_{k}$ in an action space $\mathcal{A}$, based on the observable state $s_{k}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{k}=\mathcal{R}(s_{k},a_{k})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as a function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 3.
Figure 3: DDPG scheme.
The actor-network is represented by $\mu(\cdot)$, and it determines the best action $a_{k}$ for the perimeter controller based on the current state $s_{k}$. The critic network $Q(\cdot)$ is responsible for evaluating whether a specific state-action pair at time step $k$ yields the maximal possible discounted future reward $Q(s_{k},a_{k})$. A common technique used in DDPG is to create a target actor network $\mu^{\prime}(\cdot)$ and a target critic network $Q^{\prime}(\cdot)$, which are a copy of the original actor and critic network but updated posteriorly to stabilize the training process and prevent overfitting, with the target maximal discounted future reward calculated as:
$y_{k}=r_{k}+\gamma Q^{\prime}(s_{k+1},\mu^{\prime}(s_{k+1}|\theta^{\mu^{\prime}})|\theta^{Q^{\prime}})$ (5)
The critic network can then be updated by computing the temporal difference between the predicted reward as well as the target reward and minimizing the loss for a mini-batch $N$ sampled from the replay buffer:
$L=\frac{1}{N}\sum_{k}(y_{k}-Q(s_{k},a_{k}|\theta^{Q}))^{2}$ (6)
Afterward, the actor network can be updated with the sampled deterministic policy gradient:
$I=$ $\mathbb{E}_{s_{k}\sim\rho^{\beta}}[r(s,\mu(s|\theta^{\mu}))|_{s=s_{k}}]$ (7a)
$\nabla_{\theta^{\mu}}I=$ $\mathbb{E}_{s_{k}\sim\rho^{\beta}}[\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{k},a=\mu(s_{k})}\nabla_{\theta_{\mu}}\mu(s|\theta^{\mu})|_{s=s_{k}}]$ (7b)
$\nabla_{\theta^{\mu}}I\approx$ $\frac{1}{N}\sum_{k}\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{k},a=\mu(s_{k})}\nabla_{\theta^{\mu}}\mu(s|\theta^{\mu})|_{s_{k}}$ (7c)
First developed in Horgan et al. (2018), a similar Ape-X architecture as applied in Zhou and Gayah (2021) is also adopted in this work, which allows for multiple generations of simulations during training to be included and centrally learned for the best policy. A decaying Gaussian noise is also introduced to encourage the DDPG agent to explore the state space. Starting from a certain training episode, we add disruptions into the simulation environment, such as surging traffic demand or MFD disruption. Such volatility poses additional challenges and opportunities for the learning and decision-making from the agent, leading to updates of its parameters and the capability to yield proper responses after sensing a change in a high-volatility environment.
4.3 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, proper information can be integrated into the RL algorithms to induce robustness or resilience. Following the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the benchmark work in Zhou and Gayah (2021) establishes the state as $s_{k}=[n_{ij,k},q_{ij,k}]$, with $q_{ij,k}$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij,k}$ can hardly be acquired, particularly when faced with disruptions. Therefore, we propose a state representation incorporating both the first and second order derivatives of vehicle accumulation, which can be computed as the first and second order differences from the vehicle accumulation $\Delta n_{ij,k},\Delta^{2}n_{ij,k}]$ between two consecutive time steps in the discrete form. Similar to Sirmatel and Geroliminis (2020) discussed in Section 4.1, we also tackle the data observability issue in this work, by integrating trip completion $M_{ij,k}$ and replacing OD-pair accumulation $n_{ij,k}$ with simply regional accumulation $n_{i,k}$ as well we its first and second differences. It should be noted that $M_{ij,k}$ in Sirmatel and Geroliminis (2020) refers specifically to the interregional transfer flow $M_{ij,k},(i\neq j)$. However, when considering a rather homogenous average trip length, the regional trip completion rate $M_{i}(t)$ is linearly related to the network average flow and can be approximated with ease, as illustrated in (Geroliminis and Daganzo, 2008; Zhou and Gayah, 2021). Then the intraregional completion rate can be further computed as $M_{ii,k}=M_{i,k}-M_{ij,k},(i\neq j)$. For fair comparison, we offer two sets of states for testing under either idealized full observability or real-world limited observability, which are studied in Section 6.1 and 6.2, respectively.
Idealized full observability:
$s_{k}=[n_{ij,k},\Delta n_{ij,k},\Delta^{2}n_{ij,k},M_{ij,k}]$ (8)
Real-world limited observability:
$s_{k}=[n_{i,k},\Delta n_{i,k},\Delta^{2}n_{i,k},M_{ij,k}]$ (9)
The action $a_{k}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}$ in Section 3. For the reward term $r_{k}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam,k}$ and the redundancy term $r_{red,k}$, as Eq. 4 shows. When the first and second differences $\Delta n_{ij,k},\Delta^{2}n_{ij,k}]$ or $\Delta n_{i,k},\Delta^{2}n_{i,k}]$ are incorporated into the state space $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have a minor impact on traffic performance, the operation of signal lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam,k}$ is introduced into the reward function to penalize potential oscillatory actions. While the absolute difference between two consecutive actions is always below $1$, the constant $\xi_{1}$ bounds the maximal penalty, whereas $\xi_{2}$ determines how the penalty attenuates when the change of control variables becomes small, indicating a large $\xi_{2}$ only penalizes the agent when the oscillation is substantial.
$r_{dam,k}=-\xi_{1}\sum_{i,j\in\{1,2\}}|u_{ij,k}-u_{ij,k-1}|^{\xi_{2}},\quad i\neq j\text{ and }k\neq 1$ (10)
The last term $r_{red,k}$ in the objective function $J_{anti}$ in Eq. 4 acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling (Corman et al., 2018). However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivatives in a similar way as the differences of the regional accumulation, considering real-world limited observability. Here, we summarize $r_{red,k}$ as the sum of two terms $r_{red,k}=H_{k}+\Delta H_{k}$, with $H_{k}$ being an overall term representing the first derivative and $\Delta H_{k}$ repr* esenting the second derivative, which can be further expanded as:
$H_{k}=$ $\sum_{i=1,2}H_{i,k}=\sum_{i=1,2}\omega_{h}\cdot h_{i,k}\cdot\alpha_{i,k}\cdot f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ (11a)
$\Delta H_{k}=$ $\sum_{i=1,2}\Delta H_{i,k}=\sum_{i=1,2}\omega_{\Delta h}\cdot\Delta h_{i,k}\cdot f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ (11b)
In Eq. 11, $\omega_{h}$ and $\omega_{\Delta h}$ are introduced as the weight constants for the derivatives to regulate their impact on the reward $\mathcal{R}$. $h_{i,k}$ and $\Delta h_{i,k}$ are the first and second difference of the traffic states on the MFD, $h_{i,k}$ is defined as the difference of trip completion over vehicle accumulation at the end of a time step versus at the beginning of the same time step, and the second derivative $\Delta h_{i,k}$ is calculated as the difference between the first differences of two consecutive time steps:
$h_{i,k}=\frac{M_{i,k}\ -\ M_{i,k-1}}{n_{i,k}\ -\ n_{i,k-1}}$ (12a)
$\Delta h_{k}=h_{k}-h_{k-1}$ (12b)
The binary variable $\alpha_{i,k}$ in $H_{k}$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent when accumulation decreases.
$\alpha_{i,k}=\begin{cases}1,&\text{if }n_{i,k}\geq n_{i,k-1},\\ -1,&\text{otherwise.}\end{cases}$ (13)
Here, $f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red,k}$ term when the accumulation is around the critical accumulation $n_{i,\rm crit}$, where the $r_{red,k}$ term should have the greatest impact. A modified trigonometric function is applied to realize this purpose. Other functions, such as normal distribution, should also be valid for achieving the same goal.
$f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})=\begin{cases}{\Big(1+\cos\Big(-\pi\cdot\cfrac{n_{i,\rm crit}-n_{i,k}}{n_{i,\rm crit}}\Big)\Big)}/2,&\text{if }n_{i,k}\geq n_{i,\rm crit},\\ {\Big(1+\cos\Big(-\pi\cdot\cfrac{n_{i,k}-n_{i,\rm crit}}{n_{i,\rm cap}-n_{i,\rm crit}}\Big)\Big)/2},&\text{otherwise.}\end{cases}$ (14)
We illustrate the impact of $H_{k}$ and $\Delta H_{k}$ on building redundancy on the MFD in Fig. 4 and 5. The first derivative $H_{k}$ in Fig. 4(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n_{i,k}$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n_{i,k}$ exceeds $n_{i,\rm crit}$. In Fig. 4, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.
An interesting note is that estimation uncertainty, also known as second-order uncertainty, is another factor that affects disruptions that take place unexpectedly. This is the imprecision of the learner’s current beliefs about the environment and what the antifragile terms capture. Estimation uncertainty reduces with sampling if beliefs are acquired by learning as opposed to instruction, such as anticipation through redundant overcompensation.
(a) $H_{k}$ for the first difference $h_{i,k}$.
(b) The impact on the MFD with $n_{i,k}\geq n_{i,k-1}$.
Figure 4: Illustration of the term $H_{k}$ and its effect on the MFD.
The second derivative $\Delta H_{k}$ is shown in Fig. 5. The y-axis represents how fast the traffic state changes, and the faster it reaches $n_{i,\rm crit}$, the greater the penalty will be. This observation is consistent with the redundant overcompensation and time-scale separation principles formalized in Taleb and Douady (2013) and practically applied in Axenie and Saveriano (2023). On the contrary, a reward will be given if $n$ decelerates when approaching $n_{i,\rm crit}$. Likewise, $\Delta H_{k}$ is also dependent on the normalization factor $\omega_{\Delta h}$. With $H_{k}$ and $\Delta H_{k}$, the agent learns to be conservative when regulating the perimeter control variables to reach $n_{i,\rm crit}$.
Figure 5: Illustration of the term $\Delta H_{k}$.
5 Experiment application
Following the contribution map illustrated in Fig. 1, we test our proposed antifragile perimeter control algorithm against three state-of-the-art perimeter control algorithms as benchmarks:
    MPC perimeter control proposed in Geroliminis et al. (2013).
    MPC-MHE perimeter control proposed in Sirmatel and Geroliminis (2020).
    Baseline RL perimeter control algorithm proposed in Zhou and Gayah (2021) with the original codebase [^1] https://github.com/DongqinZhou/C-D-RL.
It is worth mentioning that this work employs MPC-MHE differently in comparison to Sirmatel and Geroliminis (2020), where sensor measurement noise is assumed and addressed by MHE through state estimation. In contrast, we presume the presence of disruption but perfect sensor measurements as the real states. As a result, since there is no need to estimate the system states, MPC-MHE does not provide additional benefits over MPC alone in the context of demand disruptions. For supply disruptions, however, we introduce a disruption magnitude reduction coefficient as an additional state variable, and MPC-MHE can help estimate the coefficient and thereby the model shift, supposing such variation is well-defined and can be modeled.
Since transportation systems inherently manage the imbalance between demand and supply, real-world disruptions can also be broadly categorized as either a demand or supply disruption. A demand disruption can be easily understood as a surging traffic demand due to social events or other occurrences. Population growth and urbanization can also be considered a long-term and gradually incremental demand disruption. On the other hand, a supply disruption can represent a link-level or network-level capacity drop due to adverse weather or major accidents, with the network capacity drop being reflected by the decrease of the MFD profile, as in Lu et al. (2024). In this paper, the performance of the studied perimeter control algorithms is also tested under demand or supply disruptions, as illustrated in Fig. 6. Increasing traffic demand often leads to more frequent and severe traffic accidents (Dickerson et al., 2000). In addition to the deterministically linearly increasing magnitude of disruptions, as represented by the solid linear red line, we also tested the algorithms with uncertainties taken into account, as illustrated by the zigzag dotted red line, which will be explained in detail in the following section.
Figure 6: Illustration of simulation scenarios and their evaluation.
5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 2(b) shows. These MFDs follow Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, the lack of differentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by merely $3\%$, so that the whole MFD can be both continuous and differentiable. Other essential parameters, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 7(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.2,0.4,0.1,0.3]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,G}(t)=C_{ij}\mathcal{N}_{ij}(\mu_{ij},\sigma_{ij})$ with total number of vehicles following Gaussian as $C_{ij}=[3000,10000,2000,7000]$ and the mean and variance of the distribution being $\mu_{ij}=[1800,1800,1800,1800]$ and $\sigma_{ij}=[1200,1500,900,1200]$ in seconds. Although real-world disruptions can occur in both the outer and inner regions, exerting a negative impact on overall performance, we focus on the critical scenarios where the potential of perimeter control can be better reflected. Since the inner region has a significantly smaller MFD profile compared to its outer counterpart, a surging disruptive traffic demand is presumed to be generated in the inner region. Numerically, $q_{22,G}$ is increased to simulate traffic from within the city center, reaching a total disruption demand of $12000$ vehicles at episode $75$, accounting for approximately one-third of the normal demand.
(a) Base demand profile and demand disruption.
(b) MFD and supply disruption for the inner region.
Figure 7: Demand profile with or without surging demand.
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. In this work, we adopt a proportional decrease in both maximal accumulation and completion of the MFD, as shown in Fig. 7(b), in resemblance to Lu et al. (2024), representing lane closure events due to road construction works, traffic accidents, and natural disasters. Similar to the critical demand disruption on $q_{22,G}$, critical supply disruption requires the inner MFD being compromised while the outer MFD remains intact. Note that supply disruptions can be modeled in different ways for other events. For instance, considering the performance loss due to vehicle infrastructure interaction, Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is utilized, with a smaller $\lambda$ indicating the infrastructure is operated with higher efficiency.
The total simulation duration $T$ is 3 hours, and each time step $\Delta t$ takes $180$ seconds. The third and last hour has little demand and mainly serves as the unloading process. Another constraint is the lower and upper bounds for the perimeter control variable $u_{ij,k}\in[0.1,0.9]$ as explained in Eq. 3c. The initial vehicle accumulation is set to be $n_{ij,0}=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation. Each scenario is run $25$ times since randomness is inherent in RL and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 6, a list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The same multiplier list is applied for all scenarios that are subject to disruption magnitude uncertainties. Also, it is shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through shuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers, but with different sequences.
The most important hyperparameters for both the baseline and the proposed RL algorithms are in accordance with (Zhou and Gayah, 2021) and summarized in Table 2. Note that the minimal learning rate values are relatively larger than common RL algorithms, since we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions. The number of simulations generated under the Ape-X architecture is $32$ per training episode. The coefficients $\xi_{1}$ and $\xi_{2}$ for the damping term $r_{dam,k}$ in Eq. 10 to counter the potential oscillatory actions are set to be $1$ and $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red,k}$ is set to be $0.01$ and $0.02$. Both the prediction horizon and control horizon of MPC and MPC-MHE are set to $10$.
Table 2: List of hyperparameters.
Hyperparameter Value
Replay buffer 10,000
Sample size 1,000
Action noise initial scale 0.3
Action noise linear decay 0.003
Action noise minimal scale 0.1
Batch size 256
Target network update 5
Discount factor 0.90
Actor Critic
Initial learning rate 0.004 0.008
Learning rate decay 0.98 0.98
Minimal learning rate 0.001 0.002
Epoch 2 128
5.2 Performance evaluation
To showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 12, the main performance indicator being evaluated in this work is the TTS, which is computed by adding up the number of vehicles within the network at each second of the simulation. Although the reward in the RL algorithm is defined based on trip completion with additional damping and redundancy terms, two scenarios with the same trip completion at the end of the simulation may exhibit distinct Total Time Spent (TTS), and the one with the lower value should be regarded as having demonstrated a superior overall performance.
As Sun et al. (2024) has recently proved the fragile nature of urban road networks, fully antifragile traffic control strategies may be unattainable. When the magnitude of disruptions is sufficiently large, urban road networks will inevitably become fragile due to their real-world constraints. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is relatively more antifragile than the state-of-the-art baseline algorithms. To quantify the antifragility of different algorithms, we calculate the distribution skewness based on the samples from the last $N_{episode}=25$ incremental episodes, with $\mu$ and $\sigma$ denoting their mean and the standard deviation:
$s=\frac{1}{N_{episode}}\sum_{i=1}^{N_{episode}}\left(\frac{TTS_{i}-\mu}{\sigma}\right)^{3}$ (15)
A skewness computed as 0 indicates that the system itself or the applied algorithm makes the system neither fragile nor antifragile. Negative skewness indicates the distribution has a longer or fatter left tail and thus a higher degree of concavity in the performance function, which showcases antifragility. On the other hand, a positive skewness signifies fragility. For the detailed discussion regarding skewness and antifragility, interested readers may refer to Sun et al. (2024). Therefore, in addition to demonstrating a superior performance in terms of lower TTS, the proposed method should also yield a lower skewness than the benchmarks.
6 Results
Following the simulation setup in Section 5, we first evaluate the performance and antifragile characteristics of the studied approaches under idealized conditions with full observability. Subsequently, we examine how the RL-based algorithms perform under real-world constraints with limited observability. Each scenario is evaluated over 25 simulation runs. Performance is also assessed with and without stochastic variations for both full and limited observability settings.
6.1 Performance under full observability
Under idealized full observability, all algorithms are given perfect OD information $n_{ij}(t),q_{ij}(t)]$ without disruptions. As discussed in Section 5, MPC-MHE is only applied under supply disruption, where a linearly increasing reduction factor $r$ is assumed and can be estimated through MHE.
6.1.1 Demand disruption
Demand disruption can be understood as a surging traffic demand within a short period, as illustrated in Fig. 7(a) on top of $q_{22,G}$ as a critical scenario. With Fig. LABEL:fig:_incr_d, we present the performance curves of different studied algorithms under the cordon-shaped perimeter control environment, i.e., MPC as in Geroliminis et al. (2013) in orange, baseline RL in Zhou and Gayah (2021) in gray, and the proposed antifragile RL algorithm in blue. The first 50 episodes represent the training process for the RL-based algorithms under the static demand profile in Fig. 7(a). The TTS curves drop down and reach a comparable performance as MPC, corresponding to the conclusion in Zhou and Gayah (2021). The 50 episodes are then followed by another 25 episodes with linearly increasing demand disruption, with no stochasticity assumed on top of the linearly growing magnitude. The studied algorithms demonstrate distinct capabilities of learning and adapting to disruptions and performance variance. The proposed antifragile RL algorithm exhibits both superior performance and reduced variance, indicated by its significantly narrower shaded area. Furthermore, the performance curve of the proposed method also appears less convex than the other algorithms, demonstrating its relative antifragility. Note that due to the inherent variability of RL-based algorithms, as well as the high sensitivity of the distribution skewness on such variability, which will be computed below, we apply a sliding window of 5 episodes to smooth the performance curve.
Such superior performance between episode 507550-75 under incremental demand is further quantitatively gauged with a polar plot in Fig. LABEL:fig:_incr_d_p, which shows the performance difference in TTS reduction normalized over the baseline RL algorithm. Instead of variance, the shaded blue area highlights the performance gain of our proposed method compared to the baseline RL algorithm, showing a steady increase in TTS reduction and being largely concave. The average performance gain is $9.3\%$ and reaches $21.0\%$ after 25 episodes under incremental demand disruption. On the other hand, although MPC performs relatively better when the magnitude of disruption is low, its performance quickly draws near to the baseline RL under high demand disruptions, achieving only a $4.8\%$ improvement at peak disruption.
To quantify the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. LABEL:fig:_incr_d_sk. The value of the skewness curve at the final episode reflects the antifragility characteristics across all $25$ episodes under incremental disruptions. Since the calculation of skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase. The skewness of the baseline RL in gray rises rapidly, while the skewness of the MPC algorithm progresses even faster to $0.86$. On the contrary, the skewness curve in blue of the antifragile RL-based algorithm increases more slowly and remains largely below the baseline curve, ultimately achieving the lowest final skewness at $0.46$.
Fig. LABEL:fig:_d_un shows the performance when stochasticity is introduced into the magnitude of disruptions, following a disruption shuffling scheme discussed in Section 5.1. The results display a consistent trend with the non-stochastic scenario, but with the performance curves being less smooth, as in Fig. LABEL:fig:_incr_d_p_un, suggesting that the presence of uncertainty adds complexity to the training process. Although the performance of both RL-based methods deteriorates slightly, the proposed antifragile RL algorithm still outperforms the baseline RL by an average of $8.0\%$ and achieves $18.6\%$ performance gain at the end of the simulation under incremental demand disruptions, as well as a gain of $12.9\%$ compared to MPC. The fact that the TTS difference curve is approximately convex for MPC and concave for the antifragile RL algorithm can be quantified in Fig. LABEL:fig:_incr_d_sk_un. While the final distribution skewness of the baseline and the antifragile RL algorithms is $0.70$ and $0.56$, respectively, the skewness of MPC records a value of $0.93$, indicating its high fragility under incremental demand disruptions. An interesting phenomenon worth noticing is the spike of the skewness curve of the proposed algorithm around episode 606560-65 in LABEL:fig:_incr_d_sk_un, which can also be observed in Fig. LABEL:fig:_incr_d_sk. The spike is likely due to the interplay between multiple RL hyperparameters. Particularly after episode $50$, the replay buffer contains a mixture of training data under both disrupted and undisrupted scenarios, resulting in learning from a mixed dataset and causing a moderate deviation from optimal performance.
As the major contribution of the present research is the design of the antifragile module, which can be plugged onto other developed RL algorithms, enabling such algorithms to perform superior and become antifragile when faced with disruptions of increasing levels. Therefore, most hyperparameters of the proposed algorithm with the antifragile module should largely align with the baseline algorithm. We thus implement a sensitivity analysis on the reward weights, which is essential as a methodological contribution, and learning rates, which deviates from the original work in Zhou and Gayah (2021) due to experimental design. Table 3 summarizes the TTS at the peak disruption magnitude and the maximal skewness. We tested different magnitudes of the weight parameters in the redundancy term and found $\omega_{h}=0.01$ and $\omega_{\Delta h}=0.02$ yield the best results with both the lowest TTS and skewness. And as reward shaping heavily relies on a harmony of various reward terms, the RL agent’s decision-making can also be compromised when the weight parameters are too high. For the learning rate of the actor $lr_{a}$ and the critic $lr_{c}$, we proportionally changed the initial learning rates. Results in Table 3 demonstrate that $lr_{a}=0.008$ and $lr_{c}=0.004$ offer the best performance. Smaller learning rates will diminish the learning capability of the agent, especially under increasing disruptions. When the learning rates are high, for instance, $lr_{a}=0.016$ and $lr_{c}=0.008$, although the performance and skewness are comparable or even better than the optimal learning rate, there is a likelihood of causing gradient explosion. Likewise, $lr_{a}=0.032$ and $lr_{c}=0.016$ lead to more frequent exploding updates. For training stability, $lr_{a}=0.008$ and $lr_{c}=0.004$ are selected as the final learning rates.
Table 3: Sensitivity analysis on the weight parameters and the initial learning rates.
Weight parameters Performance Skewness Learning rate Performance Skewness
$\omega_{h}=0.000,\omega_{\Delta h}=0.000$ $1.085\cdot 10^{8}$ $0.55$ $lr_{a}=0.002,lr_{c}=0.001$ $2.282\cdot 10^{8}$ $0.92$
$\omega_{h}=0.001,\omega_{\Delta h}=0.002$ $1.048\cdot 10^{8}$ $0.49$ $lr_{a}=0.004,lr_{c}=0.002$ $1.923\cdot 10^{8}$ $0.80$
$\omega_{h}=0.010,\omega_{\Delta h}=0.020$ $1.026\cdot 10^{8}$ $0.46$ $lr_{a}=0.008,lr_{c}=0.004$ $1.026\cdot 10^{8}$ $0.46$
$\omega_{h}=0.100,\omega_{\Delta h}=0.200$ $1.217\cdot 10^{8}$ $0.62$ $lr_{a}=0.016,lr_{c}=0.008$ $1.008\cdot 10^{8}$ $0.48$
Perimeter control is based on MFD and is therefore intrinsically computationally efficient. Zhou and Gayah (2021) discussed the computation costs and highlighted that although the training time for the RL agent is higher than MPC, the computation power required for the testing phase is lower. In our experiment, both RL-based algorithms and MPC require only seconds to compute the perimeter control variables. However, as the proposed antifragile RL algorithm incorporates a higher state dimension for calculating the derivatives compared to the baseline algorithm, we further study the computation time difference for training the neural networks between these two RL-based algorithms, showing a moderate $18.7\%$ increase. As the neural network for the baseline algorithm is a relatively simple two-layer network but sufficient for solving the perimeter control problem, the training can also be achieved within seconds, and the $18.7\%$ increase is essentially negligible.
6.1.2 Supply disruption
Supply disruption can be modeled as a proportional reduction in both the capacity and the maximum vehicle accumulation on the MFD profile, as represented in Fig. 7(b). As discussed in Section 5.1, only the MFD of the inner region is reduced as a critical scenario in our simulation, where the outer region remains intact. Figure LABEL:fig:_s presents the performance curves of the studied algorithms under supply disruption. In addition to MPC (orange), baseline RL (gray), and the proposed antifragile RL (blue), the MPC-MHE variant (brown) is also included. By modeling the supply disruption magnitude coefficient $r$ as an unknown system state variable, the MHE module estimates $r$ in real time to capture the extent of the MFD reduction. Fig. LABEL:fig:_incr_s shows the performance curves of the four studied algorithms under the no disruption for the first $50$ episodes, followed by linearly reduced MFD during episode 507550-75.
The performance gain is further quantitatively analyzed in Fig. LABEL:fig:_incr_s_p, which shows an average TTS reduction of $16.7\%$ from the proposed antifragile RL compared to the baseline RL algorithm under incremental supply disruption, and reaches an ultimate performance improvement of $39.4\%$. It is worth noting that the MPC-MHE curve in brown also delivers strong performance, achieving a $31.4\%$ TTS reduction at peak disruption, highlighting the effectiveness of incorporating the supply disruption magnitude coefficient $r$ to be estimated within the MHE framework. Still, MHE relies on accurately modeling both the system dynamics and the evolution of the model reduction, which is highly challenging in real-world scenarios.
The distribution skewness up to the $n$-th episode under disruption is demonstrated in Fig. LABEL:fig:_incr_s_sk. The proposed antifragile RL method exhibits noticeably lower skewness across episodes 507550-75 compared to the other approaches, with a final skewness of $0.49$, whereas the skewness of the other three methods hovers around $1$. Note that the baseline RL algorithm exhibits a plateau in skewness toward the end of the simulation, which is caused by the fact that the network has already succumbed to a gridlock. And when the network is gridlocked, TTS merely increases linearly instead of exponentially with the growing vehicle accumulation. As a result, when overreaching the disruption threshold, the skewness of TTS initially plateaus and then gradually approaches zero. Such a phenomenon is more pronounced in the following Fig. LABEL:fig:_incr_s_sk_un, indicating that the proposed algorithm is only applicable under typical disruptions, while not for handling catastrophic scenarios where the network rapidly descends into gridlock.
When supply disruption uncertainty is introduced into the simulation, as shown in Fig. LABEL:fig:_s_un, the general trend is similar to that observed under demand disruption with uncertainty. The performance curves of MPC, MPC-MHE, and the antifragile RL algorithm all deteriorate to some extent in the presence of stochasticity. Interestingly, the performance of the baseline RL algorithm appears to improve at peak disruption compared to Fig. LABEL:fig:_incr_s. Such improvement can be attributed to the onset of gridlock and its effect on supply disruptions. When stochasticity amplifies the supply disruption magnitude coefficient, its impact is minimal since the network is already gridlocked. In contrast, when stochasticity leads to an attenuated coefficient, it reduces TTS and enhances performance. Nevertheless, the performance gain of the proposed algorithm is demonstrated quantitatively in Fig. LABEL:fig:_incr_s_p_un, showing an average improvement of $14.6\%$ and reaching $27.1\%$ at the final episode, while MPC and MPC-MHE show an ultimate $6.4\%$ and $19.1\%$ performance gain. LABEL:fig:_incr_s_sk_un summarizes the skewness of the studied algorithms, with our proposed antifragile RL achieving the lowest value of $0.70$. Although the ultimate distribution skewness of the baseline RL reaches $0.79\%$, its peak skewness is $0.90\%$ at episode $71$.
6.1.3 RL agent decision-making and explainability analysis
Previous sections report performance across the full 75 training episodes for both the state-of-the-art and the proposed antifragile algorithms, and here we examine individual representative episodes to address the explainability of the two RL-based algorithms. Specifically, we present the perimeter control variables $u_{ij}$ yielded from the RL agents and the consequent vehicle accumulation $n_{i}$.
Fig. 8 highlights episodes 50 and 75, which are the final episode before any disruption is introduced and the final episode with the highest disruption magnitude, respectively. The upper plot in Fig. 8(a) shows $u_{12}$ of both the baseline RL agent in gray and the proposed algorithm in blue, with the shaded area representing one standard deviation. Note that since the outer region is much less susceptible to traffic, both agents learn $u_{21}$ fairly easily to be at the upper bound $0.9$. The peak demand occurs at $t=1800$, resulting in the lowest $u_{12}$ for the baseline algorithm in gray. However, $n_{2}$ from the baseline algorithm depicted as the dashed gray curve in the lower plot has clearly exceeded the critical accumulation $n_{2,crit}$ shown as the red dashed line, indicating the sub-optimal decision-making of $u_{12}$. Clearly, the learned policy of the baseline algorithm with state defined as $s_{k}=[n_{ij,k},q_{ij,k}]$ is heavily influenced by $q_{ij,k}$, since $u_{12}$ from the baseline is nearly symmetric to the peak demand as in Fig. 7(a). It should also be noted that, since MFD is largely a concave function, a small deviation between the proposed $n_{2}$ as the blue dashed curve and $n_{2,crit}$ will only lead to very minor performance loss, while an increasing deviation causes exponential loss.
(a) Episode 50 no disruption.
(b) Episode 75 demand disruption.
(c) Episode 75 supply disruption.
Figure 8: Representative episodes for the perimeter control variables and the vehicle accumulation under the last episode of no disruption, peak demand disruption, and peak supply disruption scenarios.
Fig. 8(b) demonstrates $u_{ij}$ and $n_{i}$ of the final episode $75$ as the demand disruption magnitude gradually intensifies over episode 507550-75. The solid and dashed horizontal red lines represent the critical accumulation for either region $n_{1,crit}$ or $n_{2,crit}$, respectively. Due to the demand disruption on top of the base demand, both algorithms reach the lower bound $0.1$ some time before $t=1800$. The proposed RL agent remains approximately at the lower bound for a longer period, whereas the baseline RL agent quickly rebounces as a reaction to the waning of peak demand after $t=1800$, resulting in exceedingly high $n_{2}$ represented by the dashed gray curve. In contrast to Fig. 8(b), Fig. 8(c) shows that the baseline $u_{12}$ in gray has not even reached the lower bound and rebounces only shortly after the peak of demand $t=1800$, while the $u_{12}$ of the proposed algorithm in blue remains at the lower bound first and then manages the accumulation in the inner region as the dahsed blue curve somewhere around $n_{1}=3000$ after $t=5400$, corresponding to the compromised maximal completion rate under the peak supply disruption in Fig. 7(b). The above analysis on the decision-making of the RL agent demonstrates that the proposed algorithm with the state definition $s_{k}=[n_{ij,k},\Delta n_{ij,k},\Delta^{2}n_{ij,k},M_{ij,k}]$ is capable of rejecting misinformation and thus adapting more effectively to disruptions than the baseline agent, leading to antifragile performance as the disruption magnitude intensifies.
6.2 Performance under real-world limited observability
As discussed in Section 4.1, obtaining real-world measurements of $n_{ij}$ and $q_{ij}$ is highly challenging. Therefore, based on Eq. 8 and 9, we evaluate the performance of the baseline and proposed antifragile RL-based algorithms under conditions of real-world limited observability. It is important to highlight that the performance of the proposed algorithm under limited observability is not directly comparable to MPC or MPC-MHE under full observability, given their fundamentally different observability requirements. Fig. LABEL:fig:_d_obs and LABEL:fig:_d_obs_un illustrate the performance of algorithms in the presence of linearly increasing demand disruptions, modeled either deterministically or with disruption magnitude uncertainties.
In Fig. LABEL:fig:_incr_d_obs, while the TTS curve of the baseline RL algorithm in light coral is approximately the same under both full and limited observability, the performance curve of the antifragile RL algorithm under limited observability in light teal falls between that of the baseline RL and the antifragile RL with full observability. This indicates that the algorithm sacrifices some performance in exchange for depending on more realistic observable measurements. The TTS difference in Fig. LABEL:fig:_incr_d_p_obs marks the performance difference between the baseline and antifragile RL-based algorithms under limited observability, taking an average of $6.1\%$ performance improvement and reaching $15.0\%$ at the end of the simulation. It should be highlighted that the performance trade-off to achieve realistic observability, i.e., the performance difference between limited (light teal) and full observability (blue) of the proposed algorithm, is an average of $4.0\%$. The distribution skewness is quantitatively illustrated in Fig. LABEL:fig:_incr_d_sk_obs. Although taking a higher skewness value at the onset of disruption, the proposed method yields a rather low ultimate skewness of $0.49$ under limited observability, lying between the skewness observed with full observability with $0.46$ and that of the baseline RL under limited observability with $0.62$.
When under demand disruption with magnitude stochasticity as shown in Fig. LABEL:fig:_incr_d_obs_un, the proposed RL algorithm suffers from slightly larger deterioration compared to the baseline RL, mainly because the performance of the baseline RL is close to network gridlock. Still, the TTS difference in Fig. LABEL:fig:_incr_d_p_obs_un showcases the superiority of the proposed algorithm with an average of $4.2\%$ performance gain and reaches $8.2\%$ at the final episode. The performance loss for the trade-off of observability is $4.3\%$. The skewness under disruption stochasticity exhibits some volatility, which shows a value of $0.61$ and is identical to that of the baseline RL under limited observability.
In Fig. LABEL:fig:_s_obs and LABEL:fig:_s_obs_un, the performance of both algorithms under supply disruptions with limited observability follows similar trends to those under demand disruption. The shaded area in Fig. LABEL:fig:_incr_d_p_obs_un exhibits the TTS difference between the two algorithms with limited observability, which yields an average performance gain of $11.1\%$ and achieves a final value of $30.7\%$. The observability trade-off is $5.5\%$. Likewise, its skewness $0.70$ at episode $25$ is also positioned between the baseline RL with limited observability and the antifragile RL with full observability. When under disruption magnitude stochasticity in Fig. LABEL:fig:_incr_s_obs_un, as explained, the performance of the baseline RL method improves at high disruption magnitude, because the network is close to gridlock. Fig. LABEL:fig:_incr_s_p_obs_un exhibits an ultimate performance gain of $12.5\%$. Although this scenario shows the highest observability trade-off of $10.1\%$, the skewness of the proposed antifragile RL is rather low with a value of $0.69$ compared to the other algorithms.
6.3 Simulation study based on real-world data
To demonstrate the applicability of the proposed algorithm in real-world cases and test its robustness, particularly under a time-varying demand profile, we further conduct a numerical simulation using data and models derived from real-world measurements in Fig. 9. Instead of a cordon-shaped network, here we follow the network as in Fulton et al. (2025), where the city center of Zurich is divided into the left and right regions, represented by blue and red links for motorized vehicles, as shown in Fig. 9(a). Such an experimental setup avoids the complexity of through traffic passing by the city center in real-world scenarios. The green dots represent the coordinates of 500 loop detectors spread across the studied region. The measurements from the relevant detectors are used to approximate the MFD for each region. Note that due to the insignificant congestion branch in Zurich, we substituted the third-degree polynomial approximation originally applied in Zhou and Gayah (2021) with the infrastructural potential $\lambda$ method introduced in (Ambühl et al., 2020), which is a semi-analytical method that guarantees the existence of a second root, as illustrated in Fig. 9(b). The switch of the approximation function can also be used to demonstrate the insensitivity of our proposed algorithm to different MFD functional forms. The $\lambda$ method requires certain parameters with physical meanings, including free-flow speed, maximal capacity, backward wave speed, and maximal density, which are also provided in (Ambühl et al., 2020) for the studied region. In addition, by taking the average trip length and total lane length into consideration, the qkq-k MFD is then converted to the mnm-n MFD, as described in Sun et al. (2024).
(a) The city center of Zurich.
(b) Approximated MFDs with loop detector data.
Figure 9: Simulation environment constructed from the real-world network and measurements.
A realistic morning peak hour demand profile in terms of OD pairs $q_{ij,k}$ is also provided in Fulton et al. (2025) for the network described above, which contains strong time-varying stochasticity. It should be noted that real-world OD data is near-impossible to acquire, and the demand profile is calibrated through the methodology described in Ni et al. (2025). Due to the low level of inter-regional traffic flow, to ensure that perimeter control takes effect, a portion of intra-regional traffic within the right region is relocated with its new origin from the left region and destination to the right region. As both the supply and the demand based on real-world datasets in the city center of Zurich have been constructed, to demonstrate the antifragile properties of the proposed algorithm, realistic disruptions need to be approximated through historical data as well. First, the annual average traffic from the relevant loop detector measurements in 2020 is computed as the starting threshold and used as the initial training without disruptions. Then we search for the traffic flow from the beginning of 2021 that exceeds the previous threshold and update the threshold with the new peak. After searching for the whole year, a list of 10 events is summarized and illustrated by the dashed dark red curve with nodes in Fig. 10(a), with the maximal disruption magnitude as $24.5\%$ additional demand compared to the average traffic of the previous year. To tackle the time-varying demand dynamics, a moving window of 10 steps is leveraged to compute the first and second differences in Eq. 8 as well as the redundancy term in Eq. 11. The performance curves from MPC, baseline RL, and the antifragile RL algorithm are presented in Fig. 10(a), where the proposed algorithm not only converges to a lower TTS compared to the baseline algorithm, but also yields an increasingly superior performance with growing disruptions calibrated with real-world measurements. Likewise, Fig. 10(b) illustrates the distribution skewness of the three examined algorithms over the last 10 episodes and showcases that the proposed algorithm offers the lowest skewness in general, indicating its antifragile properties and the applicability under real-world scenarios.
(a) Performance curves in TTS.
(b) Skewness up to the $n$-th episode.
Figure 10: Performance curves and skewness under real-world disruptions.
6.4 Antifragile module on recent RL algorithms
So far, this study builds on the DDPG agent for perimeter control proposed in Zhou and Gayah (2021) and develops an antifragile perimeter control algorithm by augmenting the state space and shaping the reward function of the DDPG agent. The methodology functions as an add-on to the existing DDPG algorithm and enhances its performance under disruptions of previously unseen magnitudes. At the same time, however, recent advances in RL offer various advantages, including improved sample efficiency, training stability, and hyperparameter robustness, among others. Therefore, we apply our proposed methodology to the state-of-the-art algorithms to assess whether, and to what extent, such algorithms would benefit as well from the inclusion of derivatives and reward shaping under increasingly severe disruptive events. In addition to DDPG, we consider Twin Delayed Deep Deterministic policy gradient (TD3) and Soft Actor-Critic (SAC) as baseline algorithms. TD3 employs a twin critics mechanism and takes the minimum Q-value estimate to mitigate overestimation, incorporates target policy smoothing to avoid chasing sharp peaks of Q-value, and delays actor updates so that the policy is trained on steadier value estimates (Fujimoto et al., 2018). SAC is a stochastic, off-policy actor–critic algorithm that maximizes the expected return augmented with an entropy bonus and applies twin Q-networks with automatic temperature tuning to achieve stable and sample-efficient exploration (Haarnoja et al., 2018). Both TD3 and SAC have seen growing applications in the discipline of reliability and risk engineering in recent years, such as in Zhou et al. (2026); Ding et al. (2025); Cao et al. (2025).
(a) Performance curves in TTS.
(b) Normalized over baseline DDPG.
(c) Skewness up to the $n$-th episode.
Figure 11: Performance curves under incremental demand disruptions with advanced RL algorithms.
Fig. 11 displays the performance curve in TTS and the distribution skewness of the studied algorithms. The dashed curves mark the baseline algorithms for DDPG, TD3, and SAC with $s_{k}=[n_{ij,k},q_{ij,k}]$, while the solid curves represent the proposed antifragile variants with $s_{k}=[n_{ij,k},\Delta n_{ij,k},\Delta^{2}n_{ij,k},M_{ij,k}]$ and additional reward shaping terms as in Eq. 4. Clearly in Fig. 11(a), the six performance curves under incremental demand disruptions between episode 507550-75 fall into two regimes, dependent on whether the antifragile module has been leveraged. Fig. 11(b) further gauges the enhanced performance between the baselines and the proposed antifragile algorithms with an improvement varying between $20\%-30\%$ approximately. The distribution skewness is also illustrated in Fig. 11(c). The skewness of the solid curves is consistently lower than the dashed curves, with an average value of $0.57$ compared to $0.69$, demonstrating the antifragile properties of the proposed algorithm under disruptions of increasing magnitudes. Both the superior performance and the lower skewness have showcased the applicability of the proposed antifragile methodology regardless of the specific RL algorithm employed.
7 Conclusion
As disruptions are ubiquitous in urban transportation systems and are expected to increase continuously with population growth and ongoing urbanization, this work introduces the novel concept of antifragility into traffic control to tackle the growing trend of disruptions and the fragile nature of road transportation networks. First, from a terminology perspective, antifragility is compared with other related concepts commonly applied in transportation, including robustness, resilience, reliability, and adaptiveness. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust, resilient, or reliable systems from either the reliability engineering or traffic control perspective, we explore whether and how these properties can be induced with RL. Building on such concepts, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives, representing the change rate and the curvature of traffic states, to provide the RL algorithm with richer and reliable information under disruptive events. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions. In addition to the performance gain and antifragility, another contribution of this work is the consideration of real-world observability constraints, complementing the drawback of the current algorithms of utilizing hardly observable measurements.
We conducted comprehensive experiments to compare the proposed antifragile RL-based perimeter control approach against three other methods: MPC, MPC-MHE, and an RL-based algorithm. Two distinct manifestations of disruptions are examined, i.e., demand and supply disruptions. Uncertainties in the disruption magnitude are also considered. The results first demonstrate the effectiveness of our proposed antifragile RL-based algorithm, through its superior performance in terms of lower TTS. More importantly, by putting forward a novel method for quantifying antifragility through computing its performance distribution skewness, we confirm that the proposed algorithm exhibits greater antifragility with the lowest skewness among all the methods examined, delivering increasingly better performance as the magnitude of disruptions grows. Moreover, the performance and antifragile properties are studied with real-world limited observability, showing that a certain degree of performance can be traded for using only accessible sensor measurements in reality. We further validated the superior performance and relative antifragile properties of the proposed algorithm with a real-world case study. Realistic demand, supply, and disruptions are created from multiple sources of data gathered in the city center of Zurich. At last, since the proposed methodology can function as an add-on to other contemporary RL algorithms, several more advanced baselines, i.e., TD3 and SAC, are also involved. Results demonstrated the broad applicability of our proposed approach to other contemporary baselines.
Several limitations of the current study need to be addressed as well. For instance, this work primarily studies numerical simulation, with Section 6.3 showcasing a numerical simulation created from real-world network, demand, and disruption data. It should be acknowledged that real-world operation with the proposed antifragile RL-based algorithm can be more accurately validated through traffic microsimulation software like SUMO. However, due to the complexity of selecting one or multiple functional strings of traffic lights to form the perimeter in a complicated real-world network, such works are generally carried out as individual research, as in Kampitakis and Vlahogianni (2025), which is listed as one of the future directions for the present study. Furthermore, measurements and model uncertainties are commonly studied when designing perimeter control algorithms to validate their robustness. However, to uphold the clarity of this work, that is, studying antifragility instead of robustness, and also due to our introduction of uncertainties across the episodes with incremental disruptions, we only studied such time-varying demand uncertainties in the case study created from real-world datasets.
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation and control of transportation systems, continuously improving the performance of a network under unforeseen disruptions with learning-based algorithms. Additionally, it introduces a new paradigm for evaluating system operation under disruptive conditions through computing the distribution skewness. Moreover, the concept is well generic to be extended to other traffic control systems and potentially to control systems in general that are subject to growing disruptions.
8 CRediT authorship contribution statement
Linghang Sun: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft. Michail A. Makridis: Conceptualization, Methodology, Supervision, Writing - review & editing. Alexander Genser: Methodology, Visualization, Writing - review & editing. Cristian Axenie: Conceptualization, Project administration, Resources, Writing - review & editing. Margherita Grossi: Project administration, Resources, Writing - review & editing. Anastasios Kouvelas: Supervision, Writing - review & editing.
9 Acknowledgement
This research was primarily funded by the Huawei Munich Research Center through the Antigones project. Additional financial support has been received from the Swiss State Secretariat for Education, Research, and Innovation (SERI) with contract number 25.00120, as part of the European Union’s Horizon Europe project AntifragiCity with project ID 101203052.
10 Declaration of Competing Interest
This research was kindly funded by the Huawei Munich Research Center under the framework of the Antigones project, with one of our co-authors being employed at the said company. Otherwise, the authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Graphical illustration of concepts related to antifragility
Fig. 12 graphically illustrates the differences in robustness, resilience, adaptiveness, and progressive antifragility. In brief, robustness is about the system’s resistance to minor disturbances, whereas resilience represents the ability to recover from major disruptions. Adaptiveness enables the system to adapt to disruptions with magnitudes falling within a similar range, while the performance of a progressive antifragile system exhibits a concave response with linearly growing disruptions, either due to the inherent properties of the system or through antifragile control strategies.
Figure 12: Illustration for robustness, resilience, adaptiveness, and antifragility.